### v1

# Edwards-Wilkinson limit for a stochastic advection-diffusion PDE

Abstract.
We consider a diffusion in a Gaussian random environment that is white in time, and study the large-scale behavior of the quenched density with respect to the Lebesgue measure. We show that under diffusive rescaling, the fluctuations of the density converge to a Gaussian limit, described by an additive stochastic heat equation. In the case where the environment is divergence-free, our result can be interpreted as computing the scaling limit of the first-order correction to the quenched Central Limit Theorem.
Dipartimento di Matematica, Università di Pisa, Largo Bruno Pontecorvo 5, 56127 Pisa, Italy. sotirios.kotitsas@dm.unipi.it
SKLMS, Academy of Mathematics and Systems Science, Chinese Academy of Sciences, Beijing 100190, China; and School of Mathematical Sciences, University of Chinese Academy of Sciences, Beijing 100049, China. luodj@amss.ac.cn
Dipartimento di Matematica, Università di Pisa, Largo Bruno Pontecorvo 5, 56127 Pisa, Italy. mario.maurelli@unipi.it
1. Introduction
We are interested in the large-scale behavior of the following SPDE
(1.1) $\partial_{t}\theta+\nabla\cdot(V\circ\theta)=\kappa\Delta\theta,$
where $V(t,x)$ is a vector Gaussian noise and $\circ$ denotes Stratonovich integration. The noise is centered, white in time, and with the following correlation structure
(1.2) $\mathbb{E}[V_{i}(t,x)V_{j}(s,y)]=\delta(t-s)Q_{i,j}(x-y),$
for an appropriate function $Q:\mathbb{R}^{d}\rightarrow\mathbb{R}^{d\times d}$.
Our main interest in (1.1) stems from the fact that we can interpret the solution as the density of a diffusion in a random environment. Specifically, we can consider the following diffusion
(1.3) $\mathrm{d}X_{t}=V(t,X_{t})\,\mathrm{d}t+\sqrt{2\kappa}\,\mathrm{d}B_{t}.$
Then, for almost every realization of $V$, $\theta(t,x)$ is the density of $X_{t}$ with respect to the Lebesque measure.
Since the vector field $V(t,x)$ is white in time, it is not straightforward to give meaning to (1.3), and therefore, the previous observation is purely formal. Nevertheless, it is possible to make this rigorous by setting up a solution theory for this SDE. This is done, for example, using Kunita’s theory of stochastic flows [21], or the theory developed in [22], in the case where $V$ has a rough correlation function. See also [10] for a streamlined version of Kunita’s arguments. This solution theory gives a meaning to both (1.1) and (1.3), and makes the connection between them rigorous.
As such, studying the large-scale behavior of (1.1), yields ‘local’ information for a diffusion in a (white-in-time) random environment. This can be done in different scaling regimes. Specifically, [1] distinguishes three different scaling regimes: the diffusive regime, the moderate deviation regime, and the large deviation regime. These correspond to studying the diffusion (1.3) under a diffusive rescaling, under a moderate tilting of the diffusion, or the large deviation behavior, respectively. Moreover, in [1], the authors point out a very interesting connection to the KPZ equation and the KPZ universality class [3]. These conjectures were recently proved rigorously in [24, 8] in $d=1$, and for the moderate deviation regime [^1] The model studied in [24, 8] is a discrete analog of (1.1). However, their methods also apply to the continuous case, as is pointed out in [8], Section 6.3. We also refer to [4] for a study of a related, integrable model, under the large deviation regime.
Here, we are interested in the behavior of (1.1) in the diffusive scaling regime. In [1], it is conjectured that, in this regime, the fluctuations of $\theta(t,x)$, viewed as a random field, fall into the Edwards-Wilkinson universality class. To this end, we point out again the reference [10]. There, the authors studied the point-wise behavior of $\theta(t,x)$ and proved that
$\sup_{x\in\mathbb{R}^{d}}\mathbb{E}\big{[}|n^{d}\theta(n^{2}t,nx)-q_{t}(x)\Psi(n^{2}t,nx)|^{2}\big{]}\rightarrow 0,$
as $n\rightarrow\infty$, where $q_{t}(x)$ is the standard $d$-dimensional heat kernel with a specific diffusivity and $\Psi(t,x)$ is an appropriate space-time stationary random field (see [10, Section 3] for more details). Instead, what we are interested in is the behavior of
(1.4) $\mathcal{X}_{n}(t,x):=n^{d/2}(\theta_{n}(t,x)-\mathbb{E}[\theta_{n}(t,x)]),$
where $\theta_{n}(t,x):=n^{d}\theta(n^{2}t,nx)$, viewed as a random element of a Sobolev space with negative order. Our main result shows that $\mathcal{X}_{n}(t,x)$ converges in distribution to an explicit Gaussian limit, confirming the predictions of [1], see Theorem 1.9 for the precise statement.
Clearly $n^{d}\theta(n^{2}t,nx)$ corresponds to the density of $n^{-1}X_{n^{2}t}$ and therefore, the result of [10] corresponds to a quenched local central limit theorem for the diffusion, with a random correction due to the presence of $\Psi$. This result yields a quenched invariance principle as well, i.e., for almost all realizations of $V(t,x)$, $(n^{-1}X_{n^{2}t})_{t\in[0,T]}$ converges to a Brownian motion with an effective diffusivity $D_{\rm eff}$.
In this context, $\mathcal{X}_{n}(t,x)$ can be seen as the next order correction to this invariance principle. Formally, we can write
$\mathbb{E}[g(n^{-1}X_{n^{2}t})|V]=\mathbb{E}[g(D_{\rm eff}B_{t})]+\mathcal{R}_{n},$
where $\mathcal{R}_{n}$ denotes a (mean zero) random term that goes to 0, almost surely as $n\rightarrow 0$. Our main result is a central limit theorem for this error term (see Remark 1.11 for more details).
Finally, we point out that by taking the noise $V(t,x)$ in (1.1) to be a vector space-time white noise, the equation is a singular SPDE, and as such it does not make any sense [^2] When $V(t,x)$ has the regularity of the spacetime white noise the product $\nabla\cdot(\theta V)$ does not make sense. . Even worse, a formal computation, using the scaling properties of the space-time white noise, shows that it is scaling supercritical for all $d\geqslant 1$, which means the theories of regularity structures [20] or paracontrolled distributions [19] cannot be used to make sense of the equation. Here, the noises we consider have better regularity, but in our scaling regime, they converge to the spacetime white noise. As such, our result can also be seen as studying the behavior of a supercritical SPDE. In fact, the Gaussian fluctuations of (1.4) that we prove here are analogous to recent results regarding fluctuations of supercritical SPDEs (see [5, 18, 11, 2]).
We end this introduction by describing the structure of the rest of the paper. In Section 1.2 we set up the solution theory for (1.1) and state our assumptions, then we present our main result in Section 1.3 and briefly describe our methods. In Section 2.1 we collect estimates on the correlation functions of the model and prove some a priori estimates for (1.1). In Section 2.2 we prove our main result, while in Section 2.3 we show that our result still holds under weaker assumptions on the correlation function of the noise, if we assume that the latter is divergence-free.
1.1. Notation
For $x\in\mathbb{R}^{d}$, we write $x^{T}$ or $x^{\ast}$ as the transposition of $x$. The notation $|x|$ stands for the usual Euclidean norm while $\langle x\rangle=(1+|x|^{2})^{1/2}$. Let $\mathcal{S}$ be the space of Schwartz test functions and denote the Fourier transform of a function $f:\mathbb{R}^{d}\to\mathbb{R}^{m}$ as
$\widehat{f}(\xi)=\mathcal{F}(f)(\xi)=\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}{\rm e}^{-{\rm i}\xi\cdot x}f(x)\,\mathrm{d}x,\quad\xi\in\mathbb{R}^{d}.$
We write as usual $L^{p}_{x}=L^{p}(\mathbb{R}^{d})$ the Lebesgue spaces with norm $\|\cdot\|_{p}=\|\cdot\|_{L^{p}_{x}},\,p\geq 1$. For $\alpha\in\mathbb{R}$, the notation $H^{\alpha}_{x}=H^{\alpha}(\mathbb{R}^{d})$ stands for the usual inhomogeneous Sobolev space on $\mathbb{R}^{d}$, with the norm
$\|f\|_{H^{\alpha}_{x}}=\bigg{(}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{2\alpha}|\widehat{f}(\xi)|^{2}\,\mathrm{d}\xi\bigg{)}^{1/2},$
while $\dot{H}^{\alpha}_{x}$ denotes the homogeneous Sobolev space where the norm is defined by replacing $\langle\xi\rangle^{2\alpha}$ with $|\xi|^{2\alpha}$. We shall adopt the same notations for spaces of vector fields on $\mathbb{R}^{d}$.
Given $T>0$ and $p,q\geq 1$, we denote $L^{p}_{t}L^{q}_{x}$ for the time-dependent space $L^{p}([0,T],L^{q}(\mathbb{R}^{d}))$; similarly, $C_{t}L^{q}_{x}$ and $C_{t}H^{\alpha}_{x}$ are abbreviations of $C([0,T],L^{q}(\mathbb{R}^{d}))$ and $C([0,T],H^{\alpha}(\mathbb{R}^{d}))$, respectively. Sometimes, we replace the subscript $t$ by $T$ to stress the length of the time interval $[0,T]$. We write $a\lesssim b$ to mean that there is some unimportant constant $C>0$ such that $a\leqslant Cb$; to emphasize the dependence of $C$ on some parameters $d,\kappa$, we use the notation $a\lesssim_{d,\kappa}b$.
Further, we will denote by $q_{t}(x)=\frac{1}{(2\pi t)^{d/2}}e^{-\frac{|x|^{2}}{2t}}$ the standard $d$-dimensional heat kernel.
Finally, we will make use of the notation $x_{1:p}=(x_{1},\dots,x_{p})$ and for a function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$, we define $g^{\otimes p}:\mathbb{R}^{pd}\rightarrow\mathbb{R}$, $g^{\otimes p}(x_{1:p}):=g(x_{1})g(x_{2})\dots g(x_{p})$.
1.2. The Setup and Assumptions
As mentioned in the introduction, we are interested in the following SPDE:
$\partial_{t}\theta+\nabla\cdot(V\circ\theta)=\kappa\Delta\theta,$
where $\circ$ denotes Stratonovich integration. Before stating our main result, we first need to give a precise meaning to (1.1), and to prove that this is well-posed. To do this, we will first define the noise term $V$, and write down an appropriate representation. Then we use this representation to write (1.1) in Itô form, leading us to a natural notion of solution to (1.1), for which we can prove existence and uniqueness. We believe that the details of these three steps are standard. Nevertheless, we write them here for the convenience of the reader.
Recall the correlation function of the noise in (1.2). We consider two cases: when $Q$ is divergence-free (the incompressible case) and when $Q$ has possibly non-zero divergence (the compressible case). For the incompressible case, we assume the following
Assumption 1.1.
The covariance function $Q$ has a Fourier transform given by
(1.5) $\widehat{Q}(\xi)=g(\xi)\bigg{(}I_{d\times d}-\frac{\xi\xi^{T}}{|\xi|^{2}}\bigg{)},$
where $g(\xi)=g(|\xi|)$ is a nonnegative radial function satisfying $g\in(L^{1}\cap L^{\infty})(\mathbb{R}^{d})$. It is easy to show that $Q(0)=2\nu I_{d\times d}$ for some $\nu>0$.
Observe that the matrix appearing in the right-hand side of (1.5) is the projection to the subspace orthogonal to $\xi$, so that $Q$ is indeed divergence-free.
If we do not wish to assume that $Q$ is divergence-free, we instead put a stronger assumption (in terms of regularity).
Assumption 1.2.
The matrix $Q$ is smooth and compactly supported, such that $Q(0)=2\nu I_{d\times d}$.
Rigorously, one usually interprets $V$ as a cylindrical Wiener process. Specifically, we define the Hilbert space $\mathbb{H}$ as the completion of $C^{\infty}_{c}(\mathbb{R}\times\mathbb{R}^{d};\mathbb{R}^{d})$ under the following inner product
$\langle\mathbf{f},\mathbf{g}\rangle_{\mathbb{H}}:=\int_{\mathbb{R}}\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\mathbf{f}(t,x)\cdot Q(x-y)\mathbf{g}(t,y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}t.$
Then over a filtered probability space $(\Omega,\mathcal{A},(\mathcal{F}_{t})_{t\geqslant 0},\mathbb{P})$ one views the noise $V$, as a mean zero Gaussian process $(V(\mathbf{h}))_{h\in\mathbb{H}}$, with covariance function given by the inner product on $\mathbb{H}$.
For our purposes, however, we will need a more refined representation of the noise. As such we interpret $V$ as the (distributional) time derivative of an appropriate $\mathcal{Q}$-Wiener process on $L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, denoted by $W_{\mathcal{Q}}$. Here, the operator $\mathcal{Q}$ is given by
$\mathcal{Q}f(x)=\int_{\mathbb{R}^{d}}Q(x-y)f(y)\,\mathrm{d}y,$
where $f\in C^{\infty}(\mathbb{R}^{d};\mathbb{R}^{d})$. We refer to [7] for standard facts about Wiener processes on Hilbert spaces.
Observe that the operator $\mathcal{Q}$ acts as a Fourier multiplier. Furthermore, since $\mathcal{Q}$ is the correlation of the noise $V(t,x)$ and satisfies either Assumption 1.1 or Assumption 1.2, $\widehat{Q}(\xi)$ is positive definite for all $\xi\in\mathbb{R}^{d}$. This implies that $\mathcal{Q}^{\alpha}$ is a well-defined Fourier multiplier operator, for all $\alpha\in\mathbb{R}$. This allows us to define a Gaussian measure with covariance operator $\mathcal{Q}$ and with Cameron-Martin space the Hilbert space $\mathcal{H}:=\mathcal{Q}^{1/2}L^{2}(\mathbb{R}^{d};\mathbb{R}^{d})$, equipped with the inner product
$\langle\mathbf{f},\mathbf{g}\rangle:=\int_{\mathbb{R}^{d}}\mathcal{Q}^{-1/2}\mathbf{f}(x)\cdot\mathcal{Q}^{-1/2}\mathbf{g}(x)\,\mathrm{d}x.$
It can be proved that the space $\mathcal{H}$ consists of continuous, bounded vector fields. We refer to [16, Lemma 2.2] and the discussion below for more details. We also record the following lemma from the same paper, see Lemma $2.3$ therein [^3] Strictly speaking, in [16] the authors consider only the case where Assumption 1.1 holds, but it is easy to see that similar arguments can be used to establish the same results under Assumption 1.2.:
Lemma 1.3.
Let $\{\sigma_{k}\}_{k\in\mathbb{N}}$, be any orthonormal basis of $\mathcal{H}$, consisting of smooth vector fields. Then
$Q(x-y)=\sum_{n\in\mathbb{N}}\sigma_{k}(x)\sigma_{k}(y)^{T},$
where the series converges absolutely and uniformly on compact sets. If $Q$ is divergence-free, then $\sigma_{k}$ is also divergence-free, for all $k\in\mathbb{N}$.
Moreover, we have the following representation of the Fourier transform of $Q$.
Proposition 1.4.
Under Assumption 1.1 or 1.2, the following identity holds in the sense of distribution:
$\sum_{k\in\mathbb{N}}\widehat{\sigma}_{k}(\xi)\overline{\widehat{\sigma}_{k}(\eta)^{T}}=\widehat{Q}(\xi)\delta(\xi-\eta),\quad\xi,\eta\in\mathbb{R}^{d},$
where the overline means complex conjugate, and $\delta(\xi-\eta)=1$ if $\xi=\eta$ and 0 otherwise.
Proof.
Let $\phi,\psi\in\mathcal{S}(\mathbb{R}^{d};\mathbb{R}^{d})$. Then
$\lim_{N\to\infty}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\phi(\xi)^{T}\Bigg{(}\sum_{k=0}^{N}\widehat{\sigma}_{k}(\xi)\overline{\widehat{\sigma}_{k}(\eta)^{T}}\Bigg{)}\overline{\psi(\eta)}~\,\mathrm{d}\xi\,\mathrm{d}\eta$
$=\lim_{N\to\infty}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\phi}(x)^{T}\Bigg{(}\sum_{k=0}^{N}\sigma_{k}(x)\sigma_{k}(y)^{T}\Bigg{)}\overline{\widehat{\psi}(y)}~\,\mathrm{d}x\,\mathrm{d}y$
$=\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\phi}(x)^{T}\,Q(x-y)\,\overline{\widehat{\psi}(y)}~\,\mathrm{d}x\,\mathrm{d}y$
$=\int_{\mathbb{R}^{d}}\phi(\xi)^{T}\widehat{Q}(\xi)\frac{1}{(2\pi)^{d/2}}\int_{\mathbb{R}^{d}}\overline{\operatorname{e}^{{\rm i}y\cdot\xi}\widehat{\psi}(y)}~\,\mathrm{d}y\,\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}\phi(\xi)^{T}\widehat{Q}(\xi)\overline{\psi(\xi)}~\,\mathrm{d}\xi,$
where in the second identity we used the fact that $\sum_{k=0}^{N}\sigma_{k}(x)\sigma_{k}(y)^{T}\to Q(x-y)$ uniformly on any compact sets. ∎
With this representation of the covariance function, we can write
(1.6) $W_{\mathcal{Q}}(t,x)=\sum_{k\in\mathbb{N}}\sigma_{k}(x)B_{k}(t),$
where $(B_{k}(\cdot))_{k\in\mathbb{N}}$ is a collection of independent standard Brownian motions on $\mathbb{R}$, given by
$B_{k}(t)=\frac{\langle W_{\mathcal{Q}}(t),\mathcal{Q}^{-1/2}\sigma_{k}\rangle}{\|\sigma_{k}\|_{L^{2}}}.$
Finally, going back to the noise appearing in (1.1), we can write
$V(t,x)=\sum_{k\in\mathbb{N}}\sigma_{k}(x)\dot{B}_{k}(t),$
in the sense that
$V(\mathbf{h})=\sum_{k\in\mathbb{N}}\int_{\mathbb{R}_{+}}\biggl{(}\int_{\mathbb{R}^{d}}\sigma_{k}(x)\cdot\mathbf{h}(t,x)\,\mathrm{d}x\biggr{)}\,\mathrm{d}B_{k}(t),$
where $\mathbf{h}\in\mathbb{H}$.
The previous observations establish the first step listed at the beginning of this section. Now we move on to making sense of the Stratonovich integration.
At a formal level, the term $\nabla\cdot(V\circ\theta)$ is understood in the Stratonovich sense, namely
$\partial_{t}\theta+\nabla\cdot(V\circ\theta)=\kappa\Delta\theta,$
and can be written as Itô integral plus correction, that is
(1.7) $\partial_{t}\theta+\nabla\cdot(\theta V)=\nu\Delta\theta+\kappa\Delta\theta,$
where $\nu$ is as in Assumptions 1.1, 1.2. This leads us to the following notion of solution to (1.1) (see also [16, Definition 2.16]):
Definition 1.5.
Let $(\Omega,\mathcal{A},(\mathcal{F})_{t},\mathbb{P})$ be a given filtered probability space satisfying the standard assumptions, let $V$ be as above. Let $\theta_{0}\in L^{1}\cap L^{p}$, for some $p\in(1,\infty)$. A solution to (1.1) is an $(\mathcal{F})_{t}$-progressively measurable process $\theta:[0,T]\times\Omega\to L^{1}\cap L^{p}$, satisfying
  • (i)
    $\theta$ is weakly continuous and in $L^{\infty}([0,T];L^{1}\cap L^{p})\quad\mathbb{P}\text{-a.s.}$;
  • (ii)
    For all $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, we have
    $\langle\theta_{t},\phi\rangle=\langle\theta_{0},\phi\rangle+\int_{0}^{t}\langle\theta_{s},\nabla\phi\cdot V(\,\mathrm{d}s)\rangle+(\kappa+\nu)\int_{0}^{t}\langle\theta_{s},\Delta\phi\rangle\,\mathrm{d}s,$
    where we interpret the stochastic Itô integral as
    $\int_{0}^{t}\langle\theta_{s},\nabla\phi\cdot V(\,\mathrm{d}s)\rangle=\sum_{k}\int_{0}^{t}\langle\nabla\phi,\sigma_{k}\theta_{s}\rangle\,\mathrm{d}B^{k}_{s}.$
Note that the above Itô integral makes sense since, by Lemma 1.3,
$\sum_{k}|\langle\nabla\phi,\sigma_{k}\theta_{s}\rangle|^{2}$ $=\iint\theta_{s}(x)\nabla\phi(x)\cdot Q(x-y)\nabla\phi(y)\theta_{s}(y)\,\mathrm{d}x\,\mathrm{d}y$
$\leqslant|Q(0)|\,\|\nabla\phi\|_{L^{\infty}}^{2}\|\theta_{s}\|_{L^{1}}^{2}.$
Remark 1.6.
It can be shown, see for example [16, Appendix B], that, when the spectral intensity $g$ in (1.5) is decaying rapidly at infinity, that is, when $Q$ is sufficiently smooth, the Stratonovich formulation (1.1) makes sense and the equivalence with the Itô formulation (1.7) holds rigorously. Hence, the point (ii) in Definition 1.5 is equivalent to:
  • For all $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, the process $t\rightarrow\langle\phi,\theta_{t}\rangle$ is a semimartingale;
  • For all $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, we have
    $\langle\theta_{t},\phi\rangle=\langle\theta_{0},\phi\rangle+\lim_{n\rightarrow\infty}\sum_{k\leqslant n}\int_{0}^{t}\langle\nabla\phi,\theta_{s}\sigma_{k}\rangle\circ\,\mathrm{d}B_{k}(s)+\kappa\int_{0}^{t}\langle\theta_{s},\Delta\phi\rangle\,\mathrm{d}s.$
We have the following well-posedness result:
Theorem 1.7.
Assume that $Q$ satisfies Assumption 1.1 or Assumption 1.2. Then, for all $\theta_{0}\in L^{1}\cap L^{p}$, (1.1) has a unique solution, in the sense of Definition 1.5.
Further, in the case $Q$ satisfies Assumption 1.1, we have the following estimates
(1.8) $\quad\sup_{0\leqslant t\leqslant T}\|\theta_{t}\|_{L^{1}\cap L^{p}}\leqslant\|\theta_{0}\|_{L^{1}\cap L^{p}}$
and
(1.9) $\quad\sup_{0\leqslant t\leqslant T}\|\theta_{t}\|_{L^{2}}^{2}+2\kappa\int_{0}^{T}\|\nabla\theta_{t}\|_{L^{2}}^{2}\,\mathrm{d}t\leqslant 2\|\theta_{0}\|_{L^{2}}^{2}.$
In the case $Q$ satisfies Assumption 1.2, if the initial condition $\theta_{0}$ is in $C^{\infty}_{c}(\mathbb{R}^{d})$, then, for every $t>0$, the solution $\theta_{t}$ is also in $C^{\infty}_{c}(\mathbb{R}^{d})$.
Proof.
The first assertion is deduced from [10, Proposition 2.1], in the case where $Q$ satisfies Assumption 1.2, or from [16, Theorem 1.3] if $Q$ satisfies Assumption 1.1. In the latter case, the estimate (1.8) follows again from [16, Theorem 1.3], while (1.9) follows similarly by taking into account Remark 3.2 therein. The third assertion follows from the representation formula (2.5) in [10] and the smoothness of the associated stochastic flow (see also the comments before [10, Proposition 2.1]). ∎
If $\theta$ is a solution to (1.1), $\mathbb{P}$-a.s., $\theta$ is in $L^{\infty}_{t}(H^{-d/2-\varepsilon})$ by Sobolev embedding, hence $\Delta\theta$ is in $L^{\infty}_{t}(H^{-d/2-2-\varepsilon})$. By Proposition 1.4 (see also the proof of Proposition 2.6), we have ($\langle\xi\rangle:=(1+|\xi|^{2})^{1/2}$)
$\sum_{k}\|\nabla\cdot(\theta_{s}\sigma_{k})\|_{H^{-d/2-2}}^{2}$ $\leqslant\sum_{k}\|\theta_{s}\sigma_{k}\|_{H^{-d/2-1}}^{2}$
$=\sum_{k}\int|\widehat{\theta}_{s}\ast\widehat{\sigma}_{k}(\xi)|^{2}\langle\xi\rangle^{-d-2}\,\mathrm{d}\xi$
$=\int|\widehat{\theta}_{s}|^{2}\ast\operatorname{Tr}\widehat{Q}(\xi)\langle\xi\rangle^{-d-2}\,\mathrm{d}\xi$
$\lesssim\|\widehat{\theta}_{s}\|_{L^{\infty}}^{2}\|{\rm Tr}\,\widehat{Q}\|_{L^{1}}\int\langle\xi\rangle^{-d-2}\,\mathrm{d}\xi$
$\lesssim\|\theta_{s}\|_{L^{1}}^{2}.$
Hence $\sum_{k}\|\nabla\cdot(\theta_{s}\sigma_{k})\|_{H^{-d/2-2}}^{2}$ is in $L^{\infty}$ and so
$\int_{0}^{t}\nabla\cdot(\theta_{s}V(\,\mathrm{d}s))=\sum_{k}\int_{0}^{t}\nabla\cdot(\theta_{s}\sigma_{k})\,\mathrm{d}B_{k}(s)$
makes sense as $H^{-d/2-2}$-valued stochastic Itô integral and is in $C_{t}^{\gamma}(H^{-d/2-2})$ for every $\gamma<1/2$, $\mathbb{P}$-a.s.. Hence (1.1) holds as the following SDE on $H^{-d/2-2-\varepsilon}$:
$\theta_{t}=\theta_{0}-\int_{0}^{t}\nabla\cdot(\theta_{s}V(\,\mathrm{d}s))+(\kappa+\nu)\int_{0}^{t}\Delta\theta_{s}\,\mathrm{d}s.$
In particular, $\theta$ is in $C_{t}^{\gamma}(H^{-d/2-2-\varepsilon})$ for every $\gamma<1/2$, $\mathbb{P}$-a.s.. By interpolation with the $L^{\infty}_{t}(H^{-d/2-\varepsilon})$ bound, we get that, for every $\epsilon>0$, $\theta$ is in $C_{t}(H^{-d/2-\epsilon})$, $\mathbb{P}$-a.s.
These observations naturally lead us to the notion of a $H^{-d/2-\epsilon}$-mild solution to (1.7) (equivalently to (1.1)), which we will use extensively in our proofs. We say that $(\theta_{t})_{t\in[0,T]}$ is a $H^{-d/2-\epsilon}$-mild solution if
(1.10) $\theta_{t}=P_{t}\theta_{0}-\int_{0}^{t}P_{t-s}\nabla\cdot(\theta_{s}V(\,\mathrm{d}s)),$
a.s. in $H^{-d/2-\epsilon}$, where we interpret the Itô integral as in [7], $(P_{t})_{t\geqslant 0}$ be the heat semigroup generated by the operator $(\kappa+\nu)\Delta$. It is a standard fact that weak solutions to SPDEs are also mild solutions.
Proposition 1.8.
Let $\epsilon>0$, $\theta_{0}\in L^{1}\cap L^{2}$ and let $\theta$ be a weak solution of (1.1), as Definition 1.5, with $\theta_{0}$ as the initial data. Then $\theta$ is $H^{-d/2-\epsilon}$-mild solution to (1.7).
Proof.
The proof is similar to the proof from [7, Chapter 6]. Let $(\theta_{s})_{s\in[0,T]}$ be a weak solution, in the sense of Definition 1.5. Using a density argument one can show that for all $f\in C^{1}([0,t];\mathcal{S}(\mathbb{R}^{d}))$ we have
$\langle\theta_{t},f_{t}\rangle-\langle\theta_{0},f_{0}\rangle=\int_{0}^{t}\langle\theta_{s},\nabla f_{s}\cdot V(\,\mathrm{d}s)\rangle+\int_{0}^{t}\big{\langle}\theta_{s},\dot{f}_{s}+(\kappa+\nu)\Delta f_{s}\big{\rangle}\,\mathrm{d}s.$
We choose $f_{s}=P_{t-s}\phi$, where $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$. Since $\dot{f}_{s}+(\kappa+\nu)\Delta f_{s}=0$, we get
$\langle\theta_{t},\phi\rangle-\langle\theta_{0},P_{t}\phi\rangle=-\int_{0}^{t}\langle P_{t-s}\phi,\nabla\cdot(\theta_{s}V(\,\mathrm{d}s))\rangle.$
A straightforward adaptation of the proof of Proposition 2.6 below, shows that
$\int_{0}^{t}P_{t-s}\nabla\cdot(\theta_{s}V(\,\mathrm{d}s)),$
is in $H^{-d/2-\epsilon}$, when $\{\theta_{s}\}_{s\geqslant 0}$ is a predictable process such that
$\sup_{s\in[0,t]}\mathbb{E}\big{[}\|\theta_{s}\|_{L^{2}_{x}}^{2}\big{]}<\infty.$
The latter is true by (1.8), when $Q$ satisfies Assumption 1.1. If $Q$ satisfies Assumption 1.2 this bound is implied by Lemma 2.4.
As such, we can write
$\langle\theta_{t},\phi\rangle-\langle P_{t}\theta_{0},\phi\rangle=-\bigg{\langle}\int_{0}^{t}P_{t-s}\nabla\cdot(\theta_{s}V(\,\mathrm{d}s)),\phi\bigg{\rangle},$
where we also used the fact that $P_{t}$ is self-adjoint. This identity holds a.s. for all $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$. By the density of $C_{c}^{\infty}(\mathbb{R}^{d})$ in $H^{-d/2-\epsilon}$ (recall that $\theta$ is in $C_{t}(H^{-d/2-\epsilon})$ by our observations before the statement of the proposition), we conclude that (1.10) holds a.s. in $H^{-d/2-\epsilon}$. ∎
Finally, we point out that, in the case where $Q$ satisfies Assumption 1.2, (1.7) has a space-time stationary solution [10, Proposition 3.1], which we denote by $\Psi(t,x)$. Moreover, from [10, Corollary 3.2], $\mathbb{E}[\Psi(t,x)]=1$, and $\mathbb{E}[\Psi(t,x)^{2}]<\infty$. The correlation function of this field will appear in the statement of our main result. As such, define $V_{\rm eff}$ to be the symmetric matrix, such that
(1.11) $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}\mathbb{E}[\Psi(0,0)\Psi(0,x)]Q(z)\,\mathrm{d}z.$
Observe that, since $Q\in L^{1}$ and $\Psi(t,x)$ has a finite second moment, this integral is finite.
1.3. Main Result and Outline of the proof
To study the fluctuations of (1.1), we introduce a parameter $n\in\mathbb{N}$ (which we will send to $\infty$). First we rescale the initial condition to (1.1):
$\theta(0,x)=n^{-d}\varphi(x/n),$
where $\varphi(x)$ is a nice enough function. We consider the corresponding mild solution to (1.1), diffusively rescaled:
(1.12) $\theta_{n}(t,x)=n^{d}\theta(n^{2}t,nx).$
It is easy to see that the mean of $\theta_{n}$ solves the heat equation:
(1.13) $\partial_{t}\bar{\theta}=(\kappa+\nu)\Delta\bar{\theta},\quad\bar{\theta}(0)=\varphi.$
With these assumptions, (1.4) is written as
$\mathcal{X}_{n}(t,x):=n^{d/2}(\theta_{n}(t,x)-\bar{\theta}(t,x)).$
Our main result shows that $\mathcal{X}_{n}(t,x)$ has a Gaussian limit, as is expected by the CLT-type scaling.
Theorem 1.9.
Let $\alpha>d/2$ and $\gamma\in(0,1/2)$. Assume that $Q$ satisfies (1.2). Then for every $\varphi\in C^{\infty}_{c}(\mathbb{R}^{d})$, $\mathcal{X}_{n}(t,x)$ converges in distribution in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$ to $\mathcal{U}(t,x)$, the solution of the following additive stochastic heat equation:
(1.14) $\partial_{t}\mathcal{U}=(\kappa+\nu)\Delta\mathcal{U}+\nabla\cdot(\bar{\theta}\,V_{\rm eff}\xi),\quad\mathcal{U}(0,x)=0,$
where $\xi$ is a vector-valued space-time white noise, and $V_{\rm eff}$ as in (1.11).
In the case where $Q$ satisfies Assumption 1.1, the same are true, where the limiting equation is given by (1.14) with $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}Q(z)\,\mathrm{d}z$.
Remark 1.10.
The field $\Psi(t,x)$ is equal to $1$ iff $Q$ is divergence-free, see [10], thus the first part of the theorem covers the case where the correlation function is divergence-free. The point is that the divergence-free condition allows for the control of the SPDE, (1.1), even if the correlation function of the noise is not smooth, or compactly supported. In fact, Assumption 1.1 allows for Kraichnan-type noises, corresponding to $g(\xi)=\langle\xi\rangle^{-(d+\zeta)}$, for $\zeta\in(0,2)$. On the other hand, it seems possible Assumption 1.2 is too strong. Indeed, the results of Section 2.1 can be proved under the assumption that $Q$ is differentiable and Lipschitz, see Remark 2.3. Here, we work under Assumption 1.2 in order to use the results of [10], specifically, (1.15). We believe that (1.15) remains true under weaker assumptions on $Q$, but we do not pursue this here in order not to detract from our main result.
Remark 1.11.
Observe that we take as the initial data $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$. When $\varphi$ is nonnegative and $\|\varphi\|_{L^{1}}=1$, we can interpret this choice as starting the diffusion (1.3) with an initial condition $X_{0}\sim\varphi(x)\,\mathrm{d}x$. Then, we can write
$\langle\mathcal{X}_{n}(t,\cdot),g\rangle=n^{d/2}\big{(}\mathbb{E}_{\varphi}[g(n^{-1}X_{n^{2}t})|V]-\mathbb{E}[g(\tilde{B}_{t})]\big{)},$
Here, $\mathbb{E}_{\varphi}$ is the expectation with respect to the law of the diffusion (1.3), with $\varphi(x)\,\mathrm{d}x$ as the initial distribution and conditional on $V$, and $(\tilde{B}_{t})_{t\geqslant 0}$ denotes a Brownian motion with diffusivity $(\kappa+\nu)I_{d\times d}$, with $\varphi(x)\,\mathrm{d}x$ as an initial distribution.
The result of [10], implies that
$\mathbb{E}_{\varphi}[g(n^{-1}X_{n^{2}t})|V]-\mathbb{E}[g(\tilde{B}_{t})]\rightarrow 0,$
almost surely. This is a form of a quenched central limit theorem. Therefore, as mentioned in the introduction, Theorem 1.9 shows that the scaling limit of the first order correction to the quenched CLT is Gaussian with an explicit variance. Notably, we cannot take $\varphi$ to be a Dirac delta function centered at 0 (i.e., start $X_{t}$ from 0). We believe this to be an artifact of the proof, and similar methods can be used to extend our main results in this case as well.
Let us sketch the basic idea of the proof. We make use of the linearity of (1.1) and the white-in-time correlations of the noise to write (1.4) as a stochastic integral. In particular, (1.4) is a martingale. Therefore, to prove Theorem 1.9, we show that (1.4) is tight in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$, and that its quadratic variation converges to the quadratic variation of the martingale part of the mild solution of (1.14). This, combined with the Skorohod representation theorem, will allow us to conclude.
To prove tightness for (1.4), we rely on quantitative estimates for moments of
$\|\theta_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}.$
This is the content of Proposition 2.6, which is proved in Section 2.1. To prove these estimates, we need to control the expectation of $L^{p}$ norms of $\theta_{t}$, for $p>1$. This control is immediate in the case where the noise is divergence-free, as we can use (1.8). To obtain a similar control under Assumption 1.2, we rely on the correlation functions of (1.7). These functions satisfy a closed-form PDE that has a fundamental solution which, in turn, satisfies appropriate heat kernel bounds (see Propositions 2.1 and 2.2). Using these observations, we can obtain the required control of the $L^{p}$ moments of $\theta_{t}$, see Lemma 2.4.
Having these estimates at hand, we can prove tightness of the laws of $\{\mathcal{X}_{n}\}_{n\geq 1}$ in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$ (see Lemma 2.8 and Proposition 2.9). To calculate the limiting covariance, we make use of the pointwise limiting statistics of $\theta_{n}(t,x)$. In particular, under Assumption 1.2, we make use of the result of [10]. More specifically, for a fixed $\varepsilon>0$, we use the bound [^4] Actually, in [10] this is proved in the case where $\phi$ is a delta function, centered at 0, and for all $t\geqslant t_{0}$, for some $t_{0}>0$. A straightforward adaptation of the arguments in [10] can show (1.15) as well.
(1.15) $\sup_{x\in\mathbb{R}^{d}}\mathbb{E}\big{[}|\theta_{n}(t,x)-q_{t}\ast\varphi(x)\Psi(n^{2}t,nx)|^{2}\big{]}\lesssim_{\varepsilon}n^{-\gamma}$
for some $\gamma>0$, and all $t\geqslant\varepsilon$. Here $q_{t}$ is the standard heat kernel on $\mathbb{R}^{d}$.
2. Proofs
2.1. Correlation Functions and a priori estimates
Throughout this section, we work under Assumption 1.2. We consider (1.1), with initial data $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$. It will become apparent that to control moments of the $H^{-\alpha}_{x}$ norm of $\theta_{t}-\bar{\theta}_{t}$, we will need to control $L^{p}_{x}$ norms of $\theta(t,x)$. In particular, we seek to prove a bound of the form
(2.1) $\sup_{t\in[0,T]}\mathbb{E}\big{[}\|\theta_{t}\|_{L^{p}_{x}}^{r}\big{]}\lesssim 1.$
To prove this, we utilize the correlation functions of the model. More specifically, for any $p\in\mathbb{N}$, we define the $p$-th correlation function
(2.2) $\mathcal{S}_{p}(t,x_{1:p}):=\mathbb{E}[\theta(t,x_{1})\dots\theta(t,x_{p})],$
where we recall the notation $x_{1:p}=(x_{1},\dots,x_{p})$. As we will see, good pointwise bounds for the correlation functions imply bounds of the form (2.1) (see Proposition 2.2 and Lemma 2.4, below). As mentioned in the previous section, the advantage of dealing with the correlation functions, instead of $\|\theta\|_{L^{p}_{x}}$ directly, is that (2.2) satisfies an explicit parabolic PDE. Indeed, define the matrix
(2.3) $\mathcal{C}_{p}(x_{1:p})=\begin{pmatrix}(\kappa+\nu)I_{d}&&Q(x_{1}-x_{2})^{T}&&\dots&&Q(x_{1}-x_{p})^{T}\\ Q(x_{2}-x_{1})&&(\kappa+\nu)I_{d}&&\ldots&&Q(x_{2}-x_{p})^{T}\\ \vdots&&\vdots&&\ddots&&\vdots\\ Q(x_{p}-x_{1})&&Q(x_{p}-x_{2})&&\dots&&(\kappa+\nu)I_{d}\end{pmatrix}$.
We have the following proposition.
Proposition 2.1.
For all $p\in\mathbb{N}$, the correlation function (2.2) is a weak solution to
(2.4) $\partial_{t}\mathcal{S}_{p}(t,x_{1:p})=\operatorname{Tr}(\nabla^{2}\mathcal{C}_{p}(x_{1:p})\mathcal{S}_{p}(t,x_{1:p})),$
with $\mathcal{S}_{p}(0,x_{1:p})=\varphi^{\otimes p}(x_{1:p})$, where we recall that $\varphi$ is the initial data for (1.1). [^5] Here, we adopt the notation $g^{\otimes p}(x_{1:p}):=g(x_{1})g(x_{2})\dots g(x_{p})$, for a function $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$.
We point out that for a matrix $A$, we denote
$\operatorname{Tr}(\nabla^{2}(Af)):=\sum_{i,j=1}\partial_{i,j}^{2}(a_{i,j}f)$
Proof.
Let $g_{1},\dots,g_{p}\in C^{\infty}_{c}(\mathbb{R}^{d})$. From Definition 1.5 we have
(2.5) $\langle\theta_{t},g_{i}\rangle=\langle\varphi,g_{i}\rangle+\int_{0}^{t}\langle\theta_{s},(\kappa+\nu)\Delta g_{i}\rangle\,\mathrm{d}s-\int_{0}^{t}\langle\nabla g_{i},\theta_{s}V(\,\mathrm{d}s)\rangle,$
for all $i=1,\dots,p$ Now, for
$f(z_{1},\dots,z_{p}):=\prod_{i=1}^{p}z_{i},$
we apply Itô’s formula on $f(\langle\theta_{t},g_{1}\rangle,\dots,\langle\theta_{t},g_{p}\rangle)$. This yields
(2.6) $\mathrm{d}f(\langle\theta_{t},g_{1}\rangle,\dots,\langle\theta_{t},g_{p}\rangle)=\nabla f\cdot\,\mathrm{d}\Theta_{t}^{(p)}+\frac{1}{2}\sum_{i\neq j}\partial_{x_{i},x_{j}}f\,\mathrm{d}[\langle\theta_{t},g_{i}\rangle,\langle\theta_{t},g_{j}\rangle]_{t},$
where $\Theta^{(p)}_{t}:=(\langle\theta_{t},g_{1}\rangle,\dots,\langle\theta_{t},g_{p}\rangle)$. The first term on the right-hand side is equal to
(2.7) $\sum_{i=1}^{p}\partial_{x_{i}}f\,\mathrm{d}\langle\theta_{t},g_{i}\rangle=\sum_{i=1}^{p}\prod_{\begin{subarray}{c}j=1,\\ j\neq i\end{subarray}}^{p}\langle\theta_{t},g_{j}\rangle\bigl{(}\langle\theta_{t},(\kappa+\nu)\Delta g_{i}\rangle\,\mathrm{d}t-\,\mathrm{d}\langle\nabla g_{i},\theta_{t}V\rangle\bigr{)}.$
On the other hand, the second term on the right-hand side of (2.6) is equal to
$\sum_{i<j}\prod_{\begin{subarray}{c}m=1,\\ m\neq i,j\end{subarray}}\langle\theta_{t},g_{m}\rangle\,\mathrm{d}[\langle\theta_{t},g_{i}\rangle,\langle\theta_{t},g_{j}\rangle]_{t}$
(2.8) $=\sum_{i<j}\prod_{\begin{subarray}{c}m=1,\\ m\neq i,j\end{subarray}}\langle\theta_{t},g_{m}\rangle\biggl{(}\int_{\mathbb{R}^{2d}}\nabla g_{i}(x_{1})Q(x_{1}-x_{2})\nabla g_{j}(x_{2})\theta_{t}(x_{1})\theta_{t}(x_{2})\,\mathrm{d}x_{1:2}\biggr{)}\,\mathrm{d}t,$
where we used the fact that
$\,\mathrm{d}[\langle\nabla g_{i},\theta_{t}V(\,\mathrm{d}t)\rangle,\langle\nabla g_{j},\theta_{t}V(\,\mathrm{d}t)\rangle]_{t}$
$=\biggl{(}\int_{\mathbb{R}^{2d}}\nabla g_{i}(x_{1})Q(x_{1}-x_{2})\nabla g_{j}(x_{2})\theta_{t}(x_{1})\theta_{t}(x_{2})\,\mathrm{d}x_{1:2}\biggr{)}\,\mathrm{d}t,$
where $\partial_{x_{i}^{a}}$ denotes the partial derivative with respect to the $a$-th component of $x_{i}$.
Plugging (2.7) and (2.1) to (2.6), and then taking the expectation, shows that
$\langle\mathcal{S}_{p}(t,\cdot),G\rangle=\langle\varphi^{\otimes p},G\rangle-\int_{0}^{t}\langle\mathcal{S}_{p}(s,\cdot),\mathcal{C}_{p}\nabla^{2}G\rangle\,\mathrm{d}s,$
where $G(x_{1:p})=g_{1}(x_{1})g_{2}(x_{2})\dots g_{p}(x_{p})$. Arguing by density (as $\theta_{t}$ is smooth in space) concludes the proof. ∎
The point is that the PDE (2.4) is well-behaved, as the next proposition shows.
Proposition 2.2.
The PDE (2.4) has a fundamental solution, which we denote by $G_{p}(t,y_{1:p},x_{1:p})$. We also have the following heat kernel bound
(2.9) $G_{p}(t,y_{1:p},x_{1:p})\lesssim q^{\otimes p}_{ct}((x-y)_{1:p}),$
where we recall that $q_{t}(x)$ is the standard $d$-dimensional heat kernel and $c>0$ is a constant.
Proof.
Observe that the matrix $\mathcal{C}_{p}(x_{1:p})$ is uniformly elliptic and, from Assumption 1.2 it is also smooth. From standard results in parabolic PDEs, [14], this implies the existence of $G_{p}(t,y_{1:p},x_{1:p})$. To prove (2.9), we observe that the adjoint problem associated to (2.4):
$\partial_{t}f=\operatorname{Tr}(\mathcal{C}_{p}\nabla^{2}f),$
is a non-divergence form parabolic PDE with smooth, uniformly elliptic coefficients. Similarly to before, this PDE has a fundamental solution $\tilde{G}_{p}(t,y_{1:p},x_{1:p})$. From [9, Remark 5.12], $\tilde{G}_{p}$ satisfies the bound (2.9). By noticing that $G_{p}(t,y_{1:p},x_{1:p})=\tilde{G}_{p}(t,x_{1:p},y_{1:p})$, we conclude the proof. ∎
Remark 2.3.
This bound is used extensively to prove Lemma 2.4, Lemma 2.5 and Proposition 2.6. As such, it is one of the central ingredients of our arguments. We note that this proposition holds under weaker assumptions. Indeed [9] requires the coefficients to have finite Dini mean oscillation. On the other hand, to make sense of the equation using Kunita’s theory, one convenient assumption would be to take $Q$ to be differentiable and Lipschitz. As such, we expect that we can prove Theorem 1.9 under the latter assumption. As mentioned in Remark 1.10, we do not pursue this here to avoid re-proving statements we need from [10], which are proved there under Assumption 1.2.
With the bound (2.9), one can estimate moments of the $L^{2}$ norm of $\theta_{t}$, as the next lemma shows.
Lemma 2.4.
Let $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$, and start the equation (1.1) with initial data $\varphi$. Then for all $r\in\mathbb{N}$, we have
$\sup_{t\in[0,T]}\mathbb{E}\big{[}\|\theta_{t}\|_{L^{2}_{x}}^{2r}\big{]}^{1/2r}\lesssim\|\varphi\|_{L^{2}}.$
Proof.
Observe that the expectation in the statement of the lemma is equal to
$\int_{\mathbb{R}^{dr}}\mathbb{E}\bigg{[}\prod_{i=1}^{r}|\theta_{t}(x_{i})|^{2}\bigg{]}\,\mathrm{d}x_{1:r}.$
This motivates us to consider the correlation function $\mathcal{S}_{2r}(t,y_{1:2r})$. Using the fundamental solution of (2.4), we can write
$\mathcal{S}_{2r}(t,y_{1:2r})=\int_{\mathbb{R}^{2dr}}G_{2r}(t,y_{1:2r},z_{1:2r})\varphi^{\otimes 2r}(z_{1:2r})\,\mathrm{d}z_{1:2r}.$
We use the bound (2.9) which yields
$|\mathcal{S}_{2r}(t,y_{1:2r})|\lesssim\int_{\mathbb{R}^{2dr}}q_{ct}^{\otimes 2r}(y_{1:2r}-z_{1:2r})|\varphi^{\otimes 2r}(z_{1:2r})|\,\mathrm{d}z_{1:2r}.$
This holds for all $y_{1:2r}\in\mathbb{R}^{2dr}$. We choose
$y_{1}=y_{2}=x_{1},\quad y_{3}=y_{4}=x_{2},\dots,y_{2r-1}=y_{2r}=x_{r}.$
With this choice, we have
$\mathbb{E}\bigg{[}\prod_{i=1}^{r}|\theta_{t}(x_{i})|^{2}\bigg{]}=|\mathcal{S}_{2r}(t,y_{1:2r})|\lesssim\prod_{i=1}^{r}q_{ct}\ast|\varphi|(x_{i})^{2}.$
Integrating over $x_{1:r}$, and using the bound $\|q_{t}\ast\varphi\|_{L^{2}_{x}}\leqslant\|\varphi\|_{L^{2}_{x}}$, concludes the proof. ∎
We are going to need a bound on the correlation function of the rescaled solutions $\theta_{n}$. This is given by the following lemma.
Lemma 2.5.
Let $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$, and start equation (1.1) with $n^{-d}\varphi(x/n)$ as the initial data. Recall that $\theta_{n}(t,x)=n^{d}\theta(n^{2}t,nx)$. We have
$|\mathcal{S}^{n}_{2}(t,x,y)|=|\mathbb{E}[\theta_{n}(t,x)\theta_{n}(t,y)]|\lesssim q_{ct+1}(x)q_{ct+1}(y).$
Proof.
We can write
$\mathcal{S}_{2}(t,x,y)=\int_{\mathbb{R}^{2d}}G_{2}(t,x,y;z_{1},z_{2})n^{-d}\varphi(z_{1}/n)n^{-d}\varphi(z_{2}/n)\,\mathrm{d}z_{1}\,\mathrm{d}z_{2}.$
From (2.9), we get
$|\mathcal{S}_{2}(t,x,y)|\lesssim\int_{\mathbb{R}^{2d}}q_{ct}(x-z_{1})q_{ct}(y-z_{2})|n^{-d}\varphi(z_{1}/n)n^{-d}\varphi(z_{2}/n)|\,\mathrm{d}z_{1}\,\mathrm{d}z_{2}.$
Since $\varphi$ is compactly supported, we have $|n^{-d}\varphi(z_{1}/n)|\lesssim n^{-d}q_{1}(z_{1}/n)=q_{n^{2}}(z_{1})$. Therefore, we get the bound
$|\mathcal{S}_{2}(t,x,y)|\lesssim q_{ct+n^{2}}(x)q_{ct+n^{2}}(y).$
Taking $t\rightarrow n^{2}t$, $(x,y)\rightarrow(nx,ny)$, and multiplying both sides by $n^{2d}$ concludes the proof. ∎
Finally, we prove an estimate for the moments of negative Sobolev norms of the solution to (1.1). Similar estimates appear in [23]. The difference here is that we do not assume that the correlation function $Q$ is divergence-free. Recall that $\bar{\theta}$ denotes the solution to (1.13). We have
Proposition 2.6.
For all $\alpha\in(\frac{d}{2},\frac{d}{2}+1)$, and all $q\in\mathbb{N}$, we have
(2.10) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\theta_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}^{q}\Big{]}^{1/q}\lesssim_{d,\alpha,q,T}\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2},$
where $\varphi\in C_{c}^{\infty}(\mathbb{R}^{d})$ is the initial data of (1.1).
Remark 2.7.
As mentioned above, the working assumption in this section is Assumption 1.2. Nevertheless, a similar bound to (2.10) holds under Assumption 1.1 as well. This can be seen by following the same arguments as in the proof of (2.10) but using (1.8) in place of Lemma 2.4. In fact, more generally, under Assumption 1.1 we have
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\theta_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}^{q}\Big{]}^{1/q}\lesssim_{d,\alpha,\nu,q,T}\|\varphi\|_{p}\big{\|}\widehat{Q}\big{\|}_{p/(2-p)}^{1/2},$
for $p\in(1,2]$, $q\geqslant 1$ and $\alpha\in(\frac{d}{2},\frac{d}{2}+1)$. The proof is similar to the proof below, and since we are not going to use this bound, we skip the details.
Proof of Proposition 2.6.
Recall that $P_{t}=\operatorname{e}^{(\kappa+\nu)t\Delta},\,t\geqslant 0$ is the heat semigroup; by (1.10) and the mild form of the heat equation (1.13), we see that the difference $\theta_{t}-\bar{\theta}_{t}$ is nothing but the stochastic convolution
$Z_{t}:=-\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\,\mathrm{d}B_{k}(r).$
We have
(2.11) $\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/q}$ $=\bigg{[}\mathbb{E}\Big{\|}\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\,\mathrm{d}B_{k}(r)\Big{\|}_{\dot{H}^{-\alpha}}^{2q}\bigg{]}^{1/q}$
$\lesssim_{q}\bigg{[}\mathbb{E}\Big{(}\sum_{k}\int_{0}^{t}\big{\|}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\big{\|}_{\dot{H}^{-\alpha}}^{2}\,\mathrm{d}r\Big{)}^{q}\bigg{]}^{1/q},$
where in the second step we have used the Burkholder-Davis-Gundy inequality in the Hilbert space $\dot{H}^{-\alpha}$. Noting that
$\big{\|}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\big{\|}_{\dot{H}^{-\alpha}}^{2}$ $=\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\big{|}\mathcal{F}\big{(}P_{t-r}\nabla\cdot(\theta_{r}\sigma_{k})\big{)}(\xi)\big{|}^{2}\,\mathrm{d}\xi$
$=\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big{|}\xi\cdot\mathcal{F}(\theta_{r}\sigma_{k})(\xi)\big{|}^{2}\,\mathrm{d}\xi$
$=(2\pi)^{-d}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big{|}\xi\cdot(\widehat{\theta}_{r}\ast\widehat{\sigma}_{k})(\xi)\big{|}^{2}\,\mathrm{d}\xi,$
therefore,
$\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/q}$ $\lesssim_{d,q}\bigg{[}\mathbb{E}\Big{(}\sum_{k}\int_{0}^{t}\!\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\big{|}\xi\cdot(\widehat{\theta}_{r}\ast\widehat{\sigma}_{k})(\xi)\big{|}^{2}\,\mathrm{d}\xi\,\mathrm{d}r\Big{)}^{q}\bigg{]}^{1/q}$
$=\bigg{[}\mathbb{E}\Big{(}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\int_{0}^{t}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\sum_{k}\big{|}\xi\cdot(\widehat{\theta}_{r}\ast\widehat{\sigma}_{k})(\xi)\big{|}^{2}\,\mathrm{d}r\,\mathrm{d}\xi\Big{)}^{q}\bigg{]}^{1/q}.$
By Proposition 1.4, we see that
(2.12) $\sum_{k}|\xi\cdot(\widehat{\theta}_{r}\ast\widehat{\sigma}_{k})(\xi)|^{2}$
$=\lim_{N\to\infty}\sum_{k=0}^{N}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\theta}_{r}(\xi-\eta)~\xi^{T}\widehat{\sigma}_{k}(\eta)\overline{\widehat{\sigma}_{k}(\zeta)^{T}}\xi~\overline{\widehat{\theta}_{r}(\xi-\zeta)}\,\mathrm{d}\eta\,\mathrm{d}\zeta$
$=\int_{\mathbb{R}^{d}}\xi^{T}\widehat{Q}(\eta)\xi\,|\widehat{\theta}_{r}(\xi-\eta)|^{2}\,\mathrm{d}\eta.$
Plugging this into the inequality above, we get
$\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/q}\lesssim\bigg{[}\mathbb{E}\Big{(}\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha}\!\int_{0}^{t}\!{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\xi^{T}\big{(}|\widehat{\theta}_{r}|^{2}\ast\widehat{Q}\big{)}(\xi)\xi\,\mathrm{d}r\,\mathrm{d}\xi\Big{)}^{q}\bigg{]}^{1/q}.$
Using Young’s inequality, we get
$\big{|}\big{(}|\widehat{\theta}_{r}|^{2}\ast\widehat{Q}\big{)}(\xi)\big{|}\leqslant\big{\|}|\widehat{\theta}_{r}|^{2}\big{\|}_{1}\big{\|}\widehat{Q}\big{\|}_{\infty}=\|\widehat{\theta}_{r}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}=\|\theta_{r}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}$.
Plugging this to the previous bound yields
$\mathbb{E}\Big{[}\|Z_{t}\|_{\dot{H}^{-\alpha}_{x}}^{2q}\Big{]}^{1/q}\lesssim\bigg{[}\mathbb{E}\Big{(}\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\|\theta_{r}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}\,\mathrm{d}r\,\mathrm{d}\xi\Big{)}^{q}\bigg{]}^{1/q}.$
Now, from Minkowski’s inequality, we get
$\mathbb{E}\Big{[}\|\theta_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-\alpha}_{x}}^{2q}\Big{]}^{1/q}\lesssim\big{\|}\widehat{Q}\big{\|}_{\infty}\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\mathbb{E}\big{[}\|\theta_{r}\|_{2}^{2q}\big{]}^{1/q}\,\mathrm{d}r\,\mathrm{d}\xi.$
By Lemma 2.4, we have
$\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/q}$ $\lesssim\|\varphi\|_{L^{2}_{x}}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}\int_{0}^{t}\!\!\int_{\mathbb{R}^{d}}\!|\xi|^{-2\alpha+2}{\rm e}^{-2(\kappa+\nu)|\xi|^{2}(t-r)}\,\mathrm{d}r\,\mathrm{d}\xi$
$=\|\varphi\|_{L^{2}_{x}}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\frac{1-{\rm e}^{-2(\kappa+\nu)|\xi|^{2}t}}{2(\kappa+\nu)}\,\mathrm{d}\xi$
$\leqslant\|\varphi\|_{L^{2}_{x}}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\big{[}(2\nu)^{-1}\wedge(|\xi|^{2}t)\big{]}\,\mathrm{d}\xi.$
Using spherical coordinates, one has
$\int_{\mathbb{R}^{d}}|\xi|^{-2\alpha}\big{[}(2\nu)^{-1}\wedge(|\xi|^{2}t)\big{]}\,\mathrm{d}\xi$
$=c_{d}\int_{0}^{\infty}\rho^{-2\alpha}\big{[}(2\nu)^{-1}\wedge(\rho^{2}t)\big{]}\rho^{d-1}\,\mathrm{d}\rho$
$\lesssim_{d,\alpha}t\int_{0}^{(2\nu t)^{-1/2}}\rho^{-2\alpha+d+1}\,\,\mathrm{d}\rho+(2\nu)^{-1}\int_{(2\nu t)^{-1/2}}^{\infty}\rho^{-2\alpha+d-1}\,\mathrm{d}\rho$
$\lesssim_{d,\alpha,\nu}t^{\alpha-d/2},$
thanks to the constraint $\alpha\in\big{(}\frac{d}{2},1+\frac{d}{2}\big{)}$. Substituting this estimate into the above inequality yields
$\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{q}\big{]}^{1/q}\leqslant\big{[}\mathbb{E}\|Z_{t}\|_{\dot{H}^{-\alpha}}^{2q}\big{]}^{1/2q}\lesssim_{q,d,\alpha}\|\varphi\|_{L^{2}_{x}}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}t^{(2\alpha-d)/4}.$
It remains to show that the above estimate can be improved by inserting $\sup_{t\in[0,T]}$ in the expectation. This can be done in the same way as the end of proof of [23, Lemma 3.1]. Hence we omit the details here. ∎
2.2. Proof of Theorem 1.9: The compressible case
Here, we prove Theorem 1.9, under the Assumption 1.2. As mentioned in the introduction, the idea is the following: First, we will prove that $(\mathcal{X}_{n})_{n}$, defined in (1.4), is tight and we will characterize the law of all limiting points.
Recall $\theta_{n}(t,x)=n^{d}\theta(n^{2}t,nx)$, which satisfies the equation
(2.13) $\partial_{t}\theta_{n}+\nabla\cdot(\theta_{n}V^{n})=(\kappa+\nu)\Delta\theta_{n},$
with the rescaled initial condition $\theta(0,x)=n^{-d}\varphi(n^{-1}x)$. Therefore, the initial condition for (2.13) is
$\theta_{n}(0,x)=n^{d}\theta(0,nx)=\varphi(x)$
which is independent of $n$.
The noise $V^{n}$ in (2.13) has covariance function $Q^{n}(x)=Q(nx),\,x\in\mathbb{R}^{d}$; therefore, it has the Fourier transform
(2.14) $\widehat{Q^{n}}(\xi)=n^{-d}\widehat{Q}(n^{-1}\xi).$
Note that $V^{n}=\dot{W}^{n}$ and thus by (1.6),
$V^{n}(t,x)=nV(n^{2}t,nx)=\sum_{k=1}^{\infty}\sigma_{k}(nx)\,n\dot{B}_{k}(n^{2}t);$
as a result,
(2.15) $W^{n}(t,x)=\sum_{k=1}^{\infty}\sigma_{k}(nx)\,n^{-1}B_{k}(n^{2}t)\stackrel{{\scriptstyle\mathcal{L}}}{{=}}\sum_{k=1}^{\infty}\sigma_{k}(nx)B_{k}(t)$
by the scaling property of Brownian motions. Finally, note that $Q^{n}(0)=Q(0)=2\nu I_{d\times d}$ is independent of $n$.
By the above discussions, for $\alpha\in\big{(}\frac{d}{2},\frac{d}{2}+1\big{)}$, we can apply Proposition 2.6 to (2.13) and obtain
(2.16) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\theta_{n}(t)-\bar{\theta}(t)\|_{\dot{H}^{-\alpha}_{x}}^{q}\Big{]}^{1/q}$ $\lesssim_{d,\alpha,q,T}\|\varphi\|_{2}\big{\|}\widehat{Q^{n}}\big{\|}_{\infty}^{1/2}$
$=\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}n^{-d/2},$
where in the second step we have used the equality in (2.14). In particular, taking $\alpha=\frac{d}{2}+\varepsilon$ with $\varepsilon\in(0,1)$, we arrive at
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\theta_{n}(t)-\bar{\theta}(t)\|_{\dot{H}^{-d/2-\varepsilon}_{x}}^{q}\Big{]}^{1/q}$ $\lesssim_{d,\alpha,q,T}\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}n^{-d/2}.$
as a result,
(2.17) $\mathbb{E}\Big{[}\sup_{t\in[0,T]}\big{\|}n^{d/2}(\theta_{n}(t)-\bar{\theta}(t))\big{\|}_{\dot{H}^{-d/2-\varepsilon}_{x}}^{q}\Big{]}^{1/q}$ $\lesssim_{d,\alpha,q,T}\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}.$
Recalling (1.4), we have proved
$\mathbb{E}\Big{[}\sup_{t\in[0,T]}\|\mathcal{X}_{n}(t)\|_{\dot{H}^{-d/2-\varepsilon}_{x}}^{q}\Big{]}^{1/q}\lesssim_{d,\alpha,q}\|\varphi\|_{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1/2}.$
Next, note that $\mathcal{X}_{n}$ has the expression
(2.18) $\mathcal{X}_{n}(t)$ $=n^{d/2}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{n}(r)V^{n}(\,\mathrm{d}r))$
$=n^{d/2}\sum_{k}\int_{0}^{t}P_{t-r}\nabla\cdot(\theta_{n}(r)\sigma^{n}_{k})\,\mathrm{d}B^{n}_{k}(r),$
where $\sigma^{n}_{k}(x)=\sigma_{k}(nx)$, $B^{n}_{k}(r)=n^{-1}B_{k}(n^{2}r)$, $k\geq 1$. We remark that for any fixed $n\geq 1$, $\{B^{n}_{k}\}_{k\geq 1}$ are mutually independent. We turn to showing that $\mathcal{X}_{n}$ converges in the weak sense.
Lemma 2.8.
Let $\alpha>d/2$ and $\delta>0$ be such that $\alpha-\delta>d/2$, then for any $q\in 2\mathbb{N}$ and $0\leqslant s<t\leqslant T$, we have
$\mathbb{E}\|\mathcal{X}_{n}(t)-\mathcal{X}_{n}(s)\|_{H^{-\alpha}}^{q}\lesssim\|\varphi\|_{2}^{q}\big{\|}\widehat{Q}\big{\|}_{\infty}^{q/2}|t-s|^{\delta q/2}.$
Before moving on to the proof, we recall the following two elementary inequalities. Let $\alpha\in\mathbb{R}$, $\rho\geqslant 0$. Then
(2.19) $\|P_{t}g\|_{H^{\alpha+\rho}}\lesssim t^{-\rho/2}\|g\|_{H^{\alpha}},$
and for $\rho\in[0,2]$
(2.20) $\|(P_{t}-I)g\|_{H^{\alpha-\rho}}\lesssim t^{\rho/2}\|g\|_{H^{\alpha}},$
where $I$ denotes the identity operator.
Proof.
The computations below are similar to the proof of Proposition 2.6. By (2.18), we have
$\mathcal{X}_{n}(t)-\mathcal{X}_{n}(s)$ $=n^{d/2}\int_{0}^{s}\sum_{k}(P_{t-r}-P_{s-r})(\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r)))\,\mathrm{d}B^{n}_{k}(r)$
$\quad+n^{d/2}\int_{s}^{t}\sum_{k}P_{t-r}(\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r)))\,\mathrm{d}B^{n}_{k}(r)$
$=:I^{n}_{1}+I^{n}_{2}.$
By Burkholder’s inequality in Hilbert space $H^{-\alpha}$, (2.19) and (2.20) we have
$\mathbb{E}\|I^{n}_{1}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}n^{qd/2}\mathbb{E}\bigg{(}\int_{0}^{s}\sum_{k}\big{\|}(P_{t-s}-I)P_{s-r}(\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r)))\big{\|}_{H^{-\alpha}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim n^{qd/2}\mathbb{E}\bigg{(}\int_{0}^{s}\sum_{k}|t-s|^{\delta}\big{\|}P_{s-r}(\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r)))\big{\|}_{H^{-\alpha+\delta}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim_{\nu}(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg{(}\int_{0}^{s}\sum_{k}\frac{1}{|s-r|^{1-\varepsilon}}\big{\|}\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r))\big{\|}_{H^{-\alpha+\delta-1+\varepsilon}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg{(}\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta+\varepsilon}}^{2}\,\mathrm{d}r\bigg{)}^{q/2},$
where $\varepsilon$ is small enough and in the third step we have used semigroup property. We have
$\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $=\sum_{k}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\big{|}\mathcal{F}(\theta_{n}(r)\sigma^{n}_{k})(\xi)\big{|}^{2}\,\mathrm{d}\xi$
$=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\sum_{k}\big{|}\big{(}\widehat{\theta}_{n}(r)\ast\widehat{\sigma^{n}_{k}}\big{)}(\xi)\big{|}^{2}\,\mathrm{d}\xi.$
Similarly to the proof of (2.12), one has
$\sum_{k}\big{|}\big{(}\widehat{\theta}_{n}(r)\ast\widehat{\sigma^{n}_{k}}\big{)}(\xi)\big{|}^{2}$ $=\sum_{k}\iint_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\widehat{\theta}_{n}(r,\xi-\eta)\overline{\widehat{\theta}_{n}(r,\xi-\zeta)}\widehat{\sigma^{n}_{k}}(\eta)\cdot\overline{\widehat{\sigma^{n}_{k}}(\zeta)}\,\mathrm{d}\eta\,\mathrm{d}\zeta$
$=\int_{\mathbb{R}^{d}}\big{|}\widehat{\theta}_{n}(r,\xi-\eta)\big{|}^{2}{\rm Tr}(\widehat{Q^{n}}(\eta))\,\mathrm{d}\eta.$
Therefore,
$\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $=\frac{1}{(2\pi)^{d}}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\Big{(}\big{|}\widehat{\theta}_{n}(r)\big{|}^{2}\ast{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\Big{)}(\xi)\,\mathrm{d}\xi.$
By Young’s inequality, for any $\xi\in\mathbb{R}^{d}$,
$\Big{(}\big{|}\widehat{\theta}_{n}(r)\big{|}^{2}\ast{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\Big{)}(\xi)\leqslant\Big{\|}\big{|}\widehat{\theta}_{n}(r)\big{|}^{2}\Big{\|}_{1}\big{\|}{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\big{\|}_{\infty}\leqslant\big{\|}\widehat{\theta}_{n}(r)\big{\|}_{2}^{2}\big{\|}\widehat{Q^{n}}\big{\|}_{\infty}$.
Recall that $\widehat{Q^{n}}=n^{-d}\widehat{Q}(n^{-1}\cdot)$, thus $\big{\|}\widehat{Q^{n}}\big{\|}_{\infty}=n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}$; as a result,
$\Big{(}\big{|}\widehat{\theta}_{n}(r)\big{|}^{2}\ast{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\Big{)}(\xi)\leqslant\big{\|}\widehat{\theta}_{n}(r)\big{\|}_{2}^{2}\,n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}=n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2}.$
Substituting this estimate into the above equality, we arrive at
$\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta+\varepsilon}}^{2}$ $\leqslant n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2}\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta-\varepsilon)}\,\mathrm{d}\xi$
$\lesssim n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2},$
where we have used the fact that the integral is finite for $\alpha-\delta-\varepsilon>d/2$; this is possible by taking $\varepsilon$ small enough since $\alpha-\delta>d/2$. To sum up, we arrive at
$\mathbb{E}\|I^{n}_{1}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}(n^{d}|t-s|^{\delta})^{q/2}\mathbb{E}\bigg{(}\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim_{q}(n^{d}|t-s|^{\delta})^{q/2}\bigg{(}\int_{0}^{s}\frac{1}{|s-r|^{1-\varepsilon}}n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\mathbb{E}\big{[}\|\theta_{n}(r)\|_{2}^{q}\big{]}^{2/q}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim_{\varepsilon,T}\big{\|}\widehat{Q}\big{\|}_{\infty}^{q/2}\|\varphi\|_{2}^{q}|t-s|^{\delta q/2},$
where we have used Minkowski’s inequality and Lemma 2.4.
Next, we estimate $I^{n}_{2}$: again by Burkholder’s inequality and (2.19),
$\mathbb{E}\|I^{n}_{2}\|_{H^{-\alpha}}^{q}$ $\lesssim_{q}n^{dq/2}\mathbb{E}\bigg{(}\int_{s}^{t}\sum_{k}\frac{1}{|t-r|^{1-\delta}}\big{\|}\nabla\cdot(\sigma^{n}_{k}\theta_{n}(r))\big{\|}_{H^{-\alpha-1+\delta}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim n^{dq/2}\mathbb{E}\bigg{(}\int_{s}^{t}\sum_{k}\frac{1}{|t-r|^{1-\delta}}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta}}^{2}\,\mathrm{d}r\bigg{)}^{q/2}.$
Repeating the above calculations, we have
$\sum_{k}\big{\|}\sigma^{n}_{k}\theta_{n}(r)\big{\|}_{H^{-\alpha+\delta}}^{2}$ $\lesssim\int_{\mathbb{R}^{d}}\langle\xi\rangle^{-2(\alpha-\delta)}\Big{(}\big{|}\widehat{\theta_{n}}(r)\big{|}^{2}\ast{\rm Tr}\big{(}\widehat{Q^{n}}\big{)}\Big{)}(\xi)\,\mathrm{d}\xi$
$\lesssim_{\alpha,\delta}n^{-d}\big{\|}\widehat{Q}\big{\|}_{\infty}\|\theta_{n}(r)\|_{2}^{2};$
as a result,
$\mathbb{E}\|I^{n}_{2}\|_{H^{-\alpha}}^{q}$ $\lesssim\big{\|}\widehat{Q}\big{\|}_{\infty}^{q/2}\|\varphi\|_{2}^{q}\bigg{(}\int_{s}^{t}\frac{1}{|t-r|^{1-\delta}}\mathbb{E}\big{[}\|\theta_{n}(r)\|_{2}^{q}\big{]}^{2/q}\,\mathrm{d}r\bigg{)}^{q/2}$
$\lesssim_{\delta}\big{\|}\widehat{Q}\big{\|}_{\infty}^{q/2}\|\varphi\|_{2}^{q}|t-s|^{\delta q/2}.$
Combining the above two estimates, we finish the proof. ∎
Proposition 2.9.
For any $\alpha>d/2$ and $\gamma\in(0,1/2)$, the laws of $\{\mathcal{X}_{n}\}$ are tight in $C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})$.
Proof.
Taking $q$ big enough in the estimate of Lemma 2.8, we conclude from Kolmogorov’s modification theorem that $\{\mathcal{X}_{n}(t)\}_{0\leqslant t\leqslant T}$ has a version which is $\mathbb{P}$-a.s. $\gamma$-Hölder continuous in $H^{-\alpha}$; we still denote this version by $\mathcal{X}_{n}$.
Moreover, we have
$\sup_{n}\mathbb{E}\|\mathcal{X}_{n}\|_{C^{\gamma}_{t}H^{-\alpha}_{x}}^{q}\lesssim C<\infty,$
thus for any $\varepsilon>0$, we deduce from [16, Corollary A.5] that the laws of $\{\mathcal{X}_{n}\}_{n}$ are tight in $C^{\gamma}([0,T];H^{-\alpha-\varepsilon}_{\rm loc})$. By the arbitrariness of $\alpha>d/2$ and $\varepsilon>0$, we conclude the desired assertion. ∎
Next, we turn to characterizing the law of all limiting points. Observe that $\mathcal{X}_{n}$ satisfies the equation
$\mathrm{d}\mathcal{X}_{n}(t)=(\kappa+\nu)\Delta\mathcal{X}_{n}(t)\,\mathrm{d}t-n^{d/2}\nabla\cdot(V^{n}(\,\mathrm{d}t)\theta_{n}(t));$
we denote the martingale part by
$M_{n}(t)=n^{d/2}\int_{0}^{t}\nabla\cdot(\theta_{n}(s)V^{n}(\,\mathrm{d}s))=n^{d/2}\sum_{k}\int_{0}^{t}\nabla\cdot(\sigma^{n}_{k}\theta_{n}(s))\,\mathrm{d}B^{n}_{k}(s).$
Similarly to Lemma 2.8, for $\beta>1+d/2$ and small $\delta>0$, we can show that, for any $q\geq 1$,
$\mathbb{E}\|M_{n}(t)-M_{n}(s)\|_{H^{-\beta}}^{q}\lesssim\|\varphi\|_{2}^{q}\big{\|}\widehat{Q}\big{\|}_{\infty}^{q/2}|t-s|^{\delta q/2};$
as a result, we have the following analog of Proposition 2.9.
Proposition 2.10.
The laws of martingales $\{M_{n}\}_{n}$ are tight in $C^{\gamma}([0,T];H^{-\beta}_{\rm loc})$ for any $\beta>1+d/2$ and $\gamma\in(0,1/2)$.
We consider the quantity $\Xi_{n}=\big{(}\mathcal{X}_{n},M_{n},\theta_{n},\{B^{n}_{k}\}_{k}\big{)}$ which takes values in the product space
$\mathcal{S}=C^{\gamma}([0,T];H^{-\alpha}_{\rm loc})\times C^{\gamma}([0,T];H^{-\beta}_{\rm loc})\times C([0,T],H^{-\alpha})\times C([0,T],\mathbb{R}^{\mathbb{N}}).$
[…]
  • (a)
    $\tilde{\Xi}=\big{(}\tilde{\mathcal{X}},\tilde{M},\tilde{\theta},\{\tilde{B}^{k}\}_{k}\big{)}\stackrel{{\scriptstyle\mathcal{L}}}{{\sim}}\mu$ and $\tilde{\Xi}_{n}=\big{(}\tilde{\mathcal{X}}_{n},\tilde{M}_{n},\tilde{\theta}_{n},\{\tilde{B}^{n}_{k}\}_{k}\big{)}\stackrel{{\scriptstyle\mathcal{L}}}{{\sim}}\mu_{n}$ for any $n\geq 1$;
  • (b)
[…]
Item (a) implies that $\tilde{\theta}=\bar{\theta}$ solves the deterministic heat equation (1.13); moreover, for any $n\geq 1$, the following equation holds in the distributional sense:
$\,\mathrm{d}\tilde{\mathcal{X}}_{n}(t)$ $=(\kappa+\nu)\Delta\tilde{\mathcal{X}}_{n}(t)\,\mathrm{d}t-\,\mathrm{d}\tilde{M}_{n}(t)$
$=(\kappa+\nu)\Delta\tilde{\mathcal{X}}_{n}(t)\,\mathrm{d}t-n^{d/2}\sum_{k}\nabla\cdot(\sigma^{n}_{k}\tilde{\theta}_{n}(s))\,\mathrm{d}\tilde{B}^{n}_{k}(t).$
[…]
$\langle\tilde{\mathcal{X}}_{n}(t),\phi\rangle=(\kappa+\nu)\int_{0}^{t}\langle\tilde{\mathcal{X}}_{n}(s),\Delta\phi\rangle\,\mathrm{d}s-\langle\tilde{M}_{n}(t),\phi\rangle;$
by item (b) above, letting $n\to\infty$ yields
$\langle\tilde{\mathcal{X}}(t),\phi\rangle=(\kappa+\nu)\int_{0}^{t}\langle\tilde{\mathcal{X}}(s),\Delta\phi\rangle\,\mathrm{d}s-\langle\tilde{M}(t),\phi\rangle.$
It remains to identify the limit object $\{\tilde{M}(t)\}_{t\in[0,T]}$. In particular, we need to show that $\{\tilde{M}(t)\}_{t\in[0,T]}$ is a Gaussian martingale, with the correct quadratic variation. We have
Lemma 2.11.
The process $\{\tilde{M}_{t}\}_{t\in[0,T]}$ is an $H^{-\beta}$-valued martingale w.r.t. the filtration $\tilde{\mathcal{F}}_{t}=\sigma\big{(}\tilde{\theta}(s),\{\tilde{B}_{k}(s)\}_{k}:s\leqslant t\big{)}$.
Proof.
Let $\tilde{\mathcal{F}}^{n}_{t}=\sigma\big{(}\tilde{\theta}_{n}(s),\{\tilde{B}^{n}_{k}(s)\}_{k}:s\leqslant t\big{)}$. We know that $\tilde{M}_{n}$ is an $H^{-\beta}$-valued continuous martingale w.r.t. $\{\tilde{\mathcal{F}}^{n}_{t}\}_{t}$; thus for any $0\leqslant s<t\leqslant T$, $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, and any bounded continuous functional $G:C([0,s],H^{-\zeta})\times C([0,s],\mathbb{R}^{\mathbb{N}})\to\mathbb{R}$, we have
$\tilde{\mathbb{E}}\big{[}\big{\langle}\tilde{M}_{n}(t)-\tilde{M}_{n}(s),\phi\big{\rangle}\,G\big{(}\tilde{\theta}_{n}(\cdot),\{\tilde{B}^{n}_{k}(\cdot)\}_{k}\big{)}\big{]}=0.$
Letting $n\to\infty$ we obtain
$\tilde{\mathbb{E}}\big{[}\big{\langle}\tilde{M}(t)-\tilde{M}(s),\phi\big{\rangle}\,G\big{(}\tilde{\theta}_{\cdot},\{\tilde{B}_{k}(\cdot)\}_{k}\big{)}\big{]}=0.$
By the arbitrariness of $0\leqslant s<t\leqslant T$, $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$ and the functional $G$, we deduce that $\tilde{M}$ is an $H^{-\beta}$-valued martingale w.r.t. $\{\tilde{\mathcal{F}}_{t}\}_{t}$. ∎
To show that $\tilde{M}$ is a Gaussian process, we need to identify its quadratic variation; in particular, we need to calculate the limit of the quadratic variation of $\langle\tilde{M}_{n}(t),\phi\rangle$, where $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$. Note that by item (a) above, for any $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$, we have
$\tilde{M}_{n}(t)=n^{d/2}\sum_{k}\int_{0}^{t}\nabla\cdot(\sigma^{n}_{k}\tilde{\theta}^{n}_{s})\,\mathrm{d}\tilde{B}^{n}_{k}(s),$
where we have written $\tilde{\theta}_{n}(s)$ as $\tilde{\theta}^{n}_{s}$ to save space. We have
Lemma 2.12.
As $n\rightarrow\infty$, we have that
$\langle\tilde{M}_{n},\phi\rangle](t)\rightarrow\int_{[0,t]}\int_{\mathbb{R}^{d}}\big{(}q_{r}\ast\varphi(x)\big{)}^{2}\nabla\phi(x)^{T}V_{\rm eff}^{2}\nabla\phi(x)\,\mathrm{d}x\,\mathrm{d}r,$
in probability. In particular, the quadratic variation of $\langle\tilde{M}_{n}(t),\phi\rangle$ converges to the quadratic variation of
$\int_{0}^{t}\langle\nabla\phi,V_{\rm eff}\,\bar{\theta}_{s}\,\xi(\,\mathrm{d}s)\rangle,$
where $\xi:=(\xi_{1},\dots,\xi_{d})$, with $\xi_{i}$ being the standard space time white noise on $\mathbb{R}^{d}$.
Proof.
We calculate
$(t)$ $=n^{d}\sum_{k}\int_{0}^{t}\big{\langle}\tilde{\theta}^{n}_{s},\sigma^{n}_{k}\cdot\nabla\phi\big{\rangle}^{2}\,\mathrm{d}s$
$=n^{d}\sum_{k}\int_{0}^{t}\iint(\tilde{\theta}^{n}_{s}\nabla\phi)(x)^{\ast}\sigma^{n}_{k}(x)\otimes\sigma^{n}_{k}(y)(\tilde{\theta}^{n}_{s}\nabla\phi)(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s$
$=n^{d}\int_{0}^{t}\iint(\tilde{\theta}^{n}_{s}\nabla\phi)(x)^{\ast}Q^{n}(x-y)(\tilde{\theta}^{n}_{s}\nabla\phi)(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s,$
where $Q^{n}(z)=Q(nz),\,z\in\mathbb{R}^{d}$. Writing $Q_{n}(z)=n^{d}Q(nz)$ yields the expression
(2.21) $\langle\tilde{M}_{n},\phi\rangle](t)=\int_{0}^{t}\int_{\mathbb{R}^{2d}}\tilde{\theta}^{n}_{s}(x)\tilde{\theta}^{n}_{s}(y)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s.$
We split the time integral over the intervals $0,\varepsilon]$ and $\varepsilon,t]$, and call the corresponding terms $T_{1}$ and $T_{2}$ respectively.
First, we show that $\mathbb{E}[|T_{1}|]=O(\varepsilon)$. Indeed we have
$\mathbb{E}[|T_{1}|]\leqslant\int_{0}^{\varepsilon}\int_{\mathbb{R}^{2d}}\mathbb{E}\big{[}|\tilde{\theta}^{n}_{s}(x)\tilde{\theta}^{n}_{s}(y)|\big{]}|\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)|\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s.$
Now, we use item (a) and Lemma 2.5 to get the bound
$\mathbb{E}[|T_{1}|]\leqslant\int_{0}^{\varepsilon}\int_{\mathbb{R}^{2d}}q_{cs+1}(x)q_{cs+1}(y)|\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)|\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s=O(\varepsilon),$
where we used the fact that the integral
$\int_{\mathbb{R}^{2d}}|\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)|\,\mathrm{d}x\,\mathrm{d}y\leqslant\|Q_{n}\|_{L^{1}}\|\nabla\phi\|_{L^{2}}^{2}=\|Q\|_{L^{1}}\|\nabla\phi\|_{L^{2}}^{2}$
is finite, as $n\rightarrow\infty$.
Since $\varepsilon$ is arbitrarily small, we only need to handle the term $T_{2}$. Recall that we denote by $\Psi(t,x)$ the space-time stationary solution to (1.7), constructed in [10]. We observe that
$\mathbb{E}\bigg{[}\bigg{|}T_{2}-\int_{\varepsilon}^{t}\int_{\mathbb{R}^{2d}}\tilde{\theta}^{n}_{s}(x)q_{s}\ast\varphi(y)\Psi(n^{2}s,ny)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s\bigg{|}\bigg{]}\rightarrow 0,$
as $n\rightarrow\infty$. Here, we made use of (1.15). Let us denote by $\tilde{T}_{2}$ the second term in the above expression. It is easy to see that
$\mathbb{E}\bigg{[}\bigg{|}\tilde{T}_{2}-\int_{\varepsilon}^{t}\int_{\mathbb{R}^{2d}}q_{s}\ast\varphi(x)\Psi(n^{2}s,nx)q_{s}\ast\varphi(y)\Psi(n^{2}s,ny)$
$\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s\bigg{|}\bigg{]}\rightarrow 0.$
This shows that we need to consider
$\mathcal{T}_{2}=\int_{\varepsilon}^{t}\int_{\mathbb{R}^{2d}}\Psi_{s}^{n}(x)\Psi_{s}^{n}(y)q_{s}\ast\varphi(x)q_{s}\ast\varphi(y)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s,$
where $\Psi_{s}^{n}(x):=\Psi(n^{2}s,nx)$. To calculate the limit, we observe that
$\mathbb{E}[\mathcal{T}_{2}]=\int_{\varepsilon}^{t}\int_{\mathbb{R}^{2d}}w(n(x-y))q_{s}\ast\varphi(x)q_{s}\ast\varphi(y)\nabla\phi(x)^{\ast}Q_{n}(x-y)\nabla\phi(y)\,\mathrm{d}x\,\mathrm{d}y\,\mathrm{d}s,$
where $w$ is the spatial covariance function of $\Psi$. From [10, Corollary 3.2, Proposition 2.3] we can deduce that
$\mathbb{E}[\mathcal{T}_{2}]\rightarrow\int_{\varepsilon}^{t}\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s,$
where $V_{\rm eff}$ is defined in (1.11). Observe that this shows that $\mathbb{E}[T_{2}]$ converges to the same limit.
We want to show that
$\mathbb{E}\bigg{[}\bigg{|}\mathcal{T}_{2}-\int_{\varepsilon}^{t}\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s\bigg{|}^{2}\bigg{]}\rightarrow 0.$
To do this, it suffices to show that
(2.22) $\mathbb{E}\big{[}\mathcal{T}_{2}^{2}\big{]}\to\bigg{(}\int_{\varepsilon}^{t}\int_{\mathbb{R}^{d}}|q_{s}\ast\varphi(x)\,V_{\rm eff}\nabla\phi(x)|^{2}\,\mathrm{d}x\,\mathrm{d}s\bigg{)}^{2}\quad\mbox{as }n\to\infty.$
We can calculate the second moment as follows:
$\mathbb{E}\big{[}\mathcal{T}_{2}^{2}\big{]}$ $=\int_{[\varepsilon,t]^{2}}\int_{\mathbb{R}^{4d}}w^{(4)}_{n}(r,s,x_{1:4})q_{r}\ast\varphi(x_{1:2})q_{s}\ast\varphi(x_{3:4})$
$\times\nabla\phi(x_{1})^{\ast}Q_{n}(x_{1}-x_{2})\nabla\phi(x_{2})\nabla\phi(x_{3})^{\ast}Q_{n}(x_{3}-x_{4})\nabla\phi(x_{4})\,\mathrm{d}x_{1:4}\,\mathrm{d}r\,\mathrm{d}s,$
where
$w^{(4)}_{n}(r,s,x_{1:4})$ $=\mathbb{E}\big{[}\Psi^{n}_{r}(x_{1})\Psi^{n}_{r}(x_{2})\Psi^{n}_{s}(x_{3})\Psi^{n}_{s}(x_{4})\big{]}.$
Making change of variables $x_{1:4}\to(x_{2}+x_{1}/n,x_{2},x_{4}+x_{3}/n,x_{4})$, we have
$\mathbb{E}\big{[}\mathcal{T}_{2}^{2}\big{]}$ $=\int_{[\varepsilon,t]^{2}}\int_{\mathbb{R}^{4d}}\tilde{w}^{(4)}_{n}(r,s,x_{1:4})q_{r}\ast\varphi(x_{2}+x_{1}/n)q_{r}\ast\varphi(x_{2})q_{s}\ast\varphi(x_{4}+x_{3}/n)q_{s}\ast\varphi(x_{4})$
(2.23) $\times\nabla\phi(x_{2}+x_{1}/n)^{\ast}Q(x_{1})\nabla\phi(x_{2})\nabla\phi(x_{4}+x_{3}/n)^{\ast}Q(x_{3})\nabla\phi(x_{4})\,\mathrm{d}x_{1:4}\,\mathrm{d}r\,\mathrm{d}s,$
where now
$\tilde{w}^{(4)}_{n}(r,s,x_{1:4})=\mathbb{E}\big{[}\Psi(n^{2}r,nx_{2}+x_{1})\Psi(n^{2}r,nx_{2})\Psi(n^{2}s,nx_{4}+x_{3})\Psi(n^{2}s,nx_{4})\big{]}.$
By Lemma 2.13, proved below, and the dominated convergence theorem, taking limit $n\to\infty$ yields
$\mathbb{E}\big{[}\mathcal{T}_{2}^{2}\big{]}$ $\to\int_{[\varepsilon,t]^{2}}\int_{\mathbb{R}^{4d}}w(x_{1})w(x_{3})\big{(}q_{r}\ast\varphi(x_{2})\big{)}^{2}\big{(}q_{s}\ast\varphi(x_{4})\big{)}^{2}$
$\quad\times\nabla\phi(x_{2})^{T}Q(x_{1})\nabla\phi(x_{2})\,\nabla\phi(x_{4})^{\ast}Q(x_{3})\nabla\phi(x_{4})\,\mathrm{d}x_{1:4}\,\mathrm{d}r\,\mathrm{d}s$
$=\bigg{[}\int_{[\varepsilon,t]}\int_{\mathbb{R}^{2d}}\big{(}q_{r}\ast\varphi(x_{2})\big{)}^{2}\nabla\phi(x_{2})^{T}(wQ)(x_{1})\nabla\phi(x_{2})\,\mathrm{d}x_{1:2}\,\mathrm{d}r\bigg{]}^{2}$
$=\bigg{[}\int_{[\varepsilon,t]}\int_{\mathbb{R}^{d}}\big{(}q_{r}\ast\varphi(x_{2})\big{)}^{2}\nabla\phi(x_{2})^{T}V_{\rm eff}^{2}\nabla\phi(x_{2})\,\mathrm{d}x_{2}\,\mathrm{d}r\bigg{]}^{2},$
which proves (2.22). This concludes the proof. ∎
Observe that Lemma 2.12 implies that the limit of the martingale $\langle\tilde{M}_{n}(\cdot),\phi\rangle$ is a Gaussian process in $t$. Furthermore, from Theorem $8.2$ in [7], we can find a space time white noise $\xi:=(\xi_{1},\dots,\xi_{d})$, defined over the same probability space, such that
$\langle\tilde{M}_{n},\phi\rangle(t)=\int_{0}^{t}\langle\nabla\phi,V_{\rm eff}\,\bar{\theta}_{s}\,\xi(\,\mathrm{d}s)\rangle,$
a.s. for all $\phi\in C^{\infty}_{c}(\mathbb{R}^{d})$. Therefore, the limiting point $\tilde{\mathcal{X}}$ satisfies the equation
$\mathcal{X}_{t}(\phi)=(k+\nu)\int_{0}^{t}\mathcal{X}_{s}(\Delta\phi)ds+\int_{0}^{t}\langle\nabla\phi,\bar{\theta}_{s}\,\xi(\,\mathrm{d}s)\rangle,$
a.s., for all $\phi\in C_{c}^{\infty}(\mathbb{R}^{d})$. This characterizes the law of $\mathcal{X}$ in $C^{\gamma}([0,T],H^{-\alpha-\epsilon}_{loc})$, which is given by the law of the weak solution to the additive heat equation (1.14). This proves Theorem 1.9, under Assumption 1.2.
We conclude this section with the proof of Lemma 2.13, referenced above.
Lemma 2.13.
The function $\tilde{w}^{(4)}_{n}(r,s,x_{1:4})$ is uniformly bounded in $n\geq 1$, $r,s\in[\varepsilon,t]$ and $x_{1:4}\in\mathbb{R}^{4d}$. Moreover, for any $\varepsilon<r\neq s<t$, it holds
$\tilde{w}^{(4)}_{n}(r,s,x_{1:4})\to w(x_{1})w(x_{3})\quad\mbox{as }n\to\infty.$
Proof.
Following [10], we denote $\theta^{[M]}(t,x)$ as the solution of the stochastic transport-diffusion equation (1.1) starting at t=Mt=-M with initial data $\theta^{[M]}(-M,\cdot)\equiv 1$; then it holds (cf. [10, Corollary 3.2])
(2.24) $\lim_{M\to\infty}\mathbb{E}\big{|}\theta^{[M]}(t,x)-\Psi(t,x)\big{|}^{2}=0.$
Using Proposition 2.2, one can prove uniform estimates on higher order moments of $\theta^{[M]}(t,x)$. Indeed, for any $p\geqslant 2$ with $p\in\mathbb{N}$, defining the correlation function
$\mathcal{S}_{p}(t+M,x_{1:p})=\mathbb{E}\big{[}\theta^{[M]}(t,x_{1})\ldots\theta^{[M]}(t,x_{p})\big{]},\quad t\geq-M,$
then similarly to the proof of Lemma 2.4, using the fact $\theta^{[M]}(-M,\cdot)\equiv 1$ and Proposition 2.2, we have
$|\mathcal{S}_{p}(t+M,x_{1:p})|$ $=\bigg{|}\int_{\mathbb{R}^{dp}}G_{p}(t+M,x_{1:p},y_{1:p})\,\mathrm{d}y_{1:p}\bigg{|}$
$\lesssim\int_{\mathbb{R}^{dp}}q_{c(t+M)}^{\otimes p}((x-y)_{1:p})\,\mathrm{d}y_{1:p}=1.$
In particular, for any $p$ even and $x_{1}=\ldots=x_{p}=x\in\mathbb{R}^{d}$, we obtain the moment estimates
$\mathbb{E}\big{[}\big{|}\theta^{[M]}(t,x)\big{|}^{p}\big{]}\lesssim 1\quad\mbox{uniformly in }t\geq 0,x\in\mathbb{R}^{d}.$
Therefore, we deduce from (2.24) that
(2.25) $\lim_{M\to\infty}\mathbb{E}\big{|}\theta^{[M]}(t,x)-\Psi(t,x)\big{|}^{p}=0.$
Combining this limit with the above uniform bound, we can obtain the first assertion on $\tilde{w}^{(4)}_{n}(r,s,x_{1:4})$.
Next, assuming $\varepsilon<r<s<t$, we have, by temporal stationarity,
$\tilde{w}^{(4)}_{n}(r,s,x_{1:4})$ $=\mathbb{E}\big{[}\Psi(n^{2}(r-s),nx_{2}+x_{1})\Psi(n^{2}(r-s),nx_{2})\Psi(0,nx_{4}+x_{3})\Psi(0,nx_{4})\big{]};$
moreover, by spatial stationarity, it holds
$\mathbb{E}\big{[}\Psi(n^{2}(r-s),nx_{2}+x_{1})\Psi(n^{2}(r-s),nx_{2})\big{]}=w(x_{1}),$
$\mathbb{E}[\Psi(0,nx_{4}+x_{3})\Psi(0,nx_{4})]=w(x_{3}).$
We have
$\tilde{w}^{(4)}_{n}(r,s,x_{1:4})-w(x_{1})w(x_{3})$
$=\mathbb{E}\big{[}\big{(}\Psi(n^{2}(r-s),nx_{2}+x_{1})\Psi(n^{2}(r-s),nx_{2})-w(x_{1})\big{)}$
$\qquad\times\big{(}\Psi(0,nx_{4}+x_{3})\Psi(0,nx_{4})-w(x_{3})\big{)}\big{]}$
$=\mathbb{E}\big{[}\big{(}\Psi(n^{2}(r-s),nx_{2}+x_{1})\Psi(n^{2}(r-s),nx_{2})-w(x_{1})\big{)}$
$\qquad\times\big{(}\Psi(0,nx_{4}+x_{3})\Psi(0,nx_{4})-\theta^{[n^{2}(r-s)]}(0,nx_{4}+x_{3})\theta^{[n^{2}(r-s)]}(0,nx_{4})\big{)}\big{]}$
$+\mathbb{E}\big{[}\big{(}\Psi(n^{2}(r-s),nx_{2}+x_{1})\Psi(n^{2}(r-s),nx_{2})-w(x_{1})\big{)}$
$\qquad\times\big{(}\theta^{[n^{2}(r-s)]}(0,nx_{4}+x_{3})\theta^{[n^{2}(r-s)]}(0,nx_{4})-w(x_{3})\big{)}\big{]}.$
Note that by construction, the stationary field $\Psi(n^{2}(r-s),x)$ depends on information of the driving field $V(t,x)$ for $t<n^{2}(r-s)<0$, while $\theta^{[n^{2}(r-s)]}(0,x)$ depends on the information of $V(t,x)$ for $t\geqslant n^{2}(r-s)$; thus, they are independent. As a result, the second expectation is equal to
$\mathbb{E}\big{[}\Psi(n^{2}(r-s),nx_{2}+x_{1})\Psi(n^{2}(r-s),nx_{2})-w(x_{1})\big{]}$
$\quad\times\mathbb{E}\big{[}\theta^{[n^{2}(r-s)]}(0,nx_{4}+x_{3})\theta^{[n^{2}(r-s)]}(0,nx_{4})-w(x_{3})\big{]}=0.$
Concerning the first expectation, we have, by spatial stationarity,
$\mathbb{E}\big{|}\Psi(0,nx_{4}+x_{3})-\theta^{[n^{2}(r-s)]}(0,nx_{4}+x_{3})\big{|}^{4}$
$=\mathbb{E}\big{|}\Psi(0,0)-\theta^{[n^{2}(r-s)]}(0,0)\big{|}^{4}\to 0\quad\mbox{as }n\to\infty,$
where the last step is due to (2.25); similarly, as $n\to\infty$,
$\mathbb{E}\big{|}\Psi(0,nx_{4})-\theta^{[n^{2}(r-s)]}(0,nx_{4})\big{|}^{4}=\mathbb{E}\big{|}\Psi(0,0)-\theta^{[n^{2}(r-s)]}(0,0)\big{|}^{4}\to 0.$
Therefore, by Hölder’s inequality, one can show that the first expectation on the right-hand side of $\tilde{w}^{(4)}_{n}(r,s,x_{1:4})-w(x_{1})w(x_{3})$ also vanishes.
Finally, the case $\varepsilon<s<r<t$ can be treated in the same way; as a consequence, we deduce that, for all $r\neq s$,
$\tilde{w}^{(4)}_{n}(r,s,x_{1:4})-w(x_{1})w(x_{3})\to 0$
as $n\to\infty$. ∎
2.3. Proof of Theorem 1.9: The incompressible case
Here we prove Theorem 1.9, under Assumption 1.1; in this case, $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}Q(x)\,\mathrm{d}x$.
We follow the same steps as before. As mentioned in Remark 2.7, Proposition 2.6 holds in this case as well. As such, the first steps in the previous section can be followed verbatim. In particular, Proposition 2.9 and Proposition 2.10 hold under Assumption 1.1 as well. As such, it remains to control the martingales
$\tilde{M}_{n}(t)=n^{d/2}\sum_{k}\int_{0}^{t}\nabla\cdot(\sigma^{n}_{k}\tilde{\theta}^{n}_{s})\,\mathrm{d}\tilde{B}^{n}_{k}(s)=n^{d/2}\sum_{k}\int_{0}^{t}(\sigma^{n}_{k}\cdot\nabla\tilde{\theta}^{n}_{s})\,\mathrm{d}\tilde{B}^{n}_{k}(s),$
where we have already used Skorohod’s representation theorem, and the fact that $\sigma^{n}_{k}(x)=\sigma_{k}(nx)$ is divergence-free. We still have (2.21), which can be written in a more compact form:
(2.26) $\langle\tilde{M}_{n},\phi\rangle]_{t}=\int_{0}^{t}\big{\langle}\tilde{\theta}^{n}_{s}\nabla\phi,Q_{n}\ast(\tilde{\theta}^{n}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s.$
We wish to show that
(2.27) $\langle\tilde{M}_{n},\phi\rangle]_{t}\rightarrow\int_{[0,t]}\int_{\mathbb{R}^{d}}\big{(}q_{r}\ast\varphi(x)\big{)}^{2}\nabla\phi(x)^{T}V_{\rm eff}^{2}\nabla\phi(x)\,\mathrm{d}x\,\mathrm{d}r,$
in probability, where now, $V_{\rm eff}^{2}=\int_{\mathbb{R}^{d}}Q(x)\,\mathrm{d}x$. We split the right-hand side of (2.26) as follows:
(2.28) $\begin{aligned} {}_{t}&=\int_{0}^{t}\big{\langle}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi,Q_{n}\ast(\tilde{\theta}^{n}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s+\int_{0}^{t}\big{\langle}\bar{\theta}_{s}\nabla\phi,Q_{n}\ast(\tilde{\theta}^{n}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s\\ &=\int_{0}^{t}\big{\langle}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi,Q_{n}\ast((\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi)\big{\rangle}\,\mathrm{d}s\\ &\quad+\int_{0}^{t}\big{\langle}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi,Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s\\ &\quad+\int_{0}^{t}\big{\langle}\bar{\theta}_{s}\nabla\phi,Q_{n}\ast((\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi)\big{\rangle}\,\mathrm{d}s\\ &\quad+\int_{0}^{t}\big{\langle}\bar{\theta}_{s}\nabla\phi,Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s.\end{aligned}$
We denote the last four terms by $J_{i},\,i=1,2,3,4$. First, one has
$J_{4}=\int_{0}^{t}\big{\langle}\bar{\theta}_{s}\nabla\phi,Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s+\int_{0}^{t}\big{\langle}\bar{\theta}_{s}\nabla\phi,V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s.$
The first part of $J_{4}$ can be estimated as
$\int_{0}^{t}\big{\|}\bar{\theta}_{s}\nabla\phi\big{\|}_{2}\big{\|}Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\|}_{2}\,\mathrm{d}s$
$\leqslant\|\varphi\|_{2}\|\nabla\phi\|_{\infty}\int_{0}^{t}\big{\|}Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\|}_{2}\,\mathrm{d}s$
which vanishes as $n\to\infty$ since $Q_{n}=n^{d}Q(n\cdot)$ is similar to an approximation of identity on $\mathbb{R}^{d}$, with $\int_{\mathbb{R}^{d}}Q(x)\,\mathrm{d}x=V_{\rm eff}^{2}$.
Next, for $J_{2}$, in the same way we have
$J_{2}$ $=\int_{0}^{t}\big{\langle}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi,Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s$
$\quad+\int_{0}^{t}\big{\langle}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi,V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\rangle}\,\mathrm{d}s$
which will be denoted as $J_{2,1}$ and $J_{2,2}$. We can regard $\nabla\phi\cdot V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)$ as a test function ($\bar{\theta}_{s}$ is smooth since it solves the heat equation (1.13)), thus by item (b) above, it is clear that $J_{2,2}$ tends to 0 as $n\to\infty$. Concerning $J_{2,1}$, we have
$|J_{2,1}|$ $\leqslant\int_{0}^{t}\big{\|}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi\big{\|}_{2}\big{\|}Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\|}_{2}\,\mathrm{d}s$
$\leqslant\|\nabla\phi\|_{\infty}\int_{0}^{t}\big{(}\|\tilde{\theta}^{n}_{s}\|_{2}+\|\bar{\theta}_{s}\|_{2}\big{)}\big{\|}Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\|}_{2}\,\mathrm{d}s$
$\leqslant 2\|\nabla\phi\|_{\infty}\|\varphi\|_{2}\int_{0}^{t}\big{\|}Q_{n}\ast(\bar{\theta}_{s}\nabla\phi)-V_{\rm eff}^{2}(\bar{\theta}_{s}\nabla\phi)\big{\|}_{2}\,\mathrm{d}s$
which also vanishes as $n\to\infty$. In the same way, we can show that $J_{3}\to 0$.
Finally, to handle the term $J_{1}$, we first prove the following lemma.
Lemma 2.14.
Let $\kappa>0$, $\theta^{n}$ be solution to
$\mathrm{d}\theta^{n}+\circ\,\mathrm{d}W^{n}\cdot\nabla\theta^{n}=\kappa\Delta\theta^{n}\,\mathrm{d}t,\quad\theta^{n}_{0}=\theta_{0}$
and $\bar{\theta}$ the solution to
$\partial_{t}\bar{\theta}=(\kappa+\nu)\Delta\bar{\theta},\quad\bar{\theta}_{0}=\theta_{0}.$
Then one has
$\mathbb{E}\int_{0}^{T}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{2}^{2}\,\mathrm{d}t\lesssim_{T}\|\theta_{0}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1-\delta}n^{-d(1-\delta)},$
where $\delta=(d+2\varepsilon)/(2+d+2\varepsilon)$.
Proof.
By Proposition 2.6, for some $\varepsilon\in(0,1)$, we have
$\mathbb{E}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-d/2-\varepsilon}}^{2}\lesssim_{T}\|\theta_{0}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}n^{-d}$
for all $t\in[0,T]$. Note that $\theta^{n}$ and $\bar{\theta}$ satisfies the following energy estimates:
-$\mathbb{P}\mbox{-a.s.},\quad\|\theta^{n}_{t}\|_{2}^{2}+2\kappa\int_{0}^{t}\|\nabla\theta^{n}_{s}\|_{2}^{2}\,\mathrm{d}s$ $\leqslant 2\|\theta_{0}\|_{2}^{2},$
$\|\bar{\theta}_{t}\|_{2}^{2}+2(\kappa+\nu)\int_{0}^{t}\|\nabla\bar{\theta}_{s}\|_{2}^{2}\,\mathrm{d}s$ $=\|\theta_{0}\|_{2}^{2}.$
The first estimate is a consequence of (1.9), since $\theta^{n}$ satisfies (1.7) with a rescaled noise term. We can get the second energy estimate by integrating by parts. Combining these estimates yields
-$\mathbb{P}\mbox{-a.s.},\quad\int_{0}^{T}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{1}}^{2}\,\mathrm{d}t\lesssim_{\kappa,T}\|\theta_{0}\|_{2}^{2}.$
By interpolation, $\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{2}\leqslant\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{1}}^{\delta}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-d/2-\varepsilon}}^{1-\delta}$, where $\delta=(d+2\varepsilon)/(2+d+2\varepsilon)$. As a result, by Cauchy’s inequality,
$\mathbb{E}\int_{0}^{T}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{2}^{2}\,\mathrm{d}t$ $\leqslant\mathbb{E}\int_{0}^{T}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{1}}^{2\delta}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-d/2-\varepsilon}}^{2(1-\delta)}\,\mathrm{d}t$
$\leqslant\bigg{[}\mathbb{E}\int_{0}^{T}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{1}}^{2}\,\mathrm{d}t\bigg{]}^{\delta}\bigg{[}\mathbb{E}\int_{0}^{T}\|\theta^{n}_{t}-\bar{\theta}_{t}\|_{\dot{H}^{-d/2-\varepsilon}}^{2}\,\mathrm{d}t\bigg{]}^{1-\delta}$
$\lesssim_{T}\|\theta_{0}\|_{2}^{2}\big{\|}\widehat{Q}\big{\|}_{\infty}^{1-\delta}n^{-d(1-\delta)}.$
Now we can estimate $J_{1}$ as follows:
$\tilde{\mathbb{E}}|J_{1}|$ $\leqslant\tilde{\mathbb{E}}\int_{0}^{T}\big{\|}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi\big{\|}_{2}\big{\|}Q_{n}\ast((\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi)\big{\|}_{2}\,\mathrm{d}s$
$\leqslant\tilde{\mathbb{E}}\int_{0}^{T}\big{\|}(\tilde{\theta}^{n}_{s}-\bar{\theta}_{s})\nabla\phi\big{\|}_{2}^{2}\|Q_{n}\|_{1}\,\mathrm{d}s,$
where in the second step we have used Young’s inequality. Next, since $\|Q_{n}\|_{1}=\|Q\|_{1}$ and $\tilde{\theta}^{n}$ has the same law as $\theta^{n}$, we have
$\tilde{\mathbb{E}}|J_{1}|$ $\leqslant\|\nabla\phi\|_{\infty}^{2}\|Q\|_{1}\,\tilde{\mathbb{E}}\int_{0}^{T}\big{\|}\tilde{\theta}^{n}_{s}-\bar{\theta}_{s}\big{\|}_{2}^{2}\,\mathrm{d}s$
$=\|\nabla\phi\|_{\infty}^{2}\|Q\|_{1}\,\mathbb{E}\int_{0}^{T}\big{\|}\theta^{n}_{s}-\bar{\theta}_{s}\big{\|}_{2}^{2}\,\mathrm{d}s\to 0$
as $n\to\infty$.
This shows (2.27). We can continue in the same way as in the previous section (after the proof of Lemma 2.12) and conclude the proof of Theorem 1.9 under Assumption 1.1 as well.
Acknowledgements. The second author is grateful to the financial supports of the National Key R&D Program of China (No. 2024YFA1012300) and the National Natural Science Foundation of China (Nos. 12090010, 12090014).